### v2

# Asymptotic error distribution for stochastic Runge–Kutta methods of strong order one

Abstract.
This work gives the asymptotic error distribution of the stochastic Runge–Kutta (SRK) method of strong order 111 applied to Stratonovich-type stochastic differential equations. For dealing with the implicitness introduced in the diffusion term, we propose a framework to derive the asymptotic error distribution of diffusion-implicit or fully implicit numerical methods, which enables us to construct a fully explicit numerical method sharing the same asymptotic error distribution as the SRK method. Further, we show that the limit distribution $U(T)$ satisfies $\mathbf{E}|U(T)|^{2}\leq e^{L_{1}T}(1+\eta_{1})T^{3}$ for some $\eta_{1}\geq 0$ only depending on the coefficients of the SRK method. Thus, we infer that $\eta_{1}$ is the key parameter reflecting the growth rate of the mean-square error of the SRK method. Especially, among the SRK methods of strong order 111, those of weak order 222 correspond to $\eta_{1}=0$, sharing the unified asymptotic error distribution, and have the smallest mean-square errors after a long time. This property is also found for the case of additive noise. It seems that we are the first to give the asymptotic error distribution of fully implicit numerical methods for stochastic differential equations.
Key words and phrases:
asymptotic error distribution, convergence in distribution, stochastic Runge–Kutta method, stochastic differential equation
This work is supported by National Natural Science Foundation of China (Nos. 12201228, 12031020 and 12171047).
AMS subject classifications: 60H35, 60H10, 60B10, 60F05
1. Introduction
In this work, we study the asymptotic error distribution of stochastic Runge–Kutta (SRK) methods applied to the following Stratonovich-type stochastic differential equation (SDE):
$\begin{cases}\mathrm{d}Y(t)=f(Y(t))\mathrm{d}t+g(Y(t))\circ\mathrm{d}\mathbf{W}(t),~t\in(0,T],\\ Y(0)=Y_{0}\in\mathbb{R}^{d},\end{cases}$ (1.1)
where $\mathbf{W}(t)=(W_{1}(t),\ldots,W_{m}(t))^{\top}_{t\in[0,T]}$ is an $m$-dimensional standard Brownian motion defined on $(\Omega,\mathcal{F},\mathbf{P})$ with respect to (w.r.t.) a filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ with $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ satisfying usual conditions, and $f:\mathbb{R}^{d}\to\mathbb{R}^{d},\,g\in\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are Lipschitz continuous such that (1.1) admits a unique strong solution. Stochastic numerical methods have been playing an important role in numerically approximating SDEs and modeling the intrinsic behaviors of SDEs over the past decades, which have received extensive attention and study (cf. [12, 16]). In the theoretical study on stochastic numerical methods, the error analysis, measuring the reliability and accuracy of stochastic numerical methods, is one of the most basic and most important subject. Compared with the fruitful results on the convergence analysis of stochastic numerical methods, including the strong convergence, weak convergence and convergence in probability, the investigation on probabilistic limit theorems for errors between numerical solutions and exact solutions has been far from sufficient. In this paper, we are interested in the asymptotic error distribution of stochastic numerical methods applied to the SDE (1.1), which is a kind of generalized central limit theorem.
1.1. Research background
The asymptotic error distribution of stochastic numerical methods is the limit distribution of the normalized error between numerical solutions and exact solutions as the step-size tends to zero, where the normalized error is the error weighted by the corresponding strong convergence order of numerical methods. The asymptotic error distribution provides the optimal strong convergence order for stochastic numerical methods and characterizes the evolution pattern of distribution of the error in the small step-size regime. It also finds some applications in the error structure of stochastic numerical methods (cf. [1]). The study on the asymptotic error distribution dated back to the work [13] for SDEs with bounded coefficients, where the authors gave a sufficient condition for Euler method to admit the asymptotic error distribution. Subsequent research by [10] provided a sufficient and necessary condition for the asymptotic error distribution of the Euler method when applied to SDEs with linearly growing coefficients. Further, [19] established the asymptotic error distribution of the Euler method for SDEs with locally Lipschitz coefficients. Recently, the asymptotic error distribution of $\theta$-method, applied to SDEs with additive noise, was obtained in [7]. We also refer to the asymptotic error distribution of numerical methods applied to stochastic integral equations ([4, 18]), Mckean–Vlasov SDEs ([14]) and SDEs driven by fractional Brownian motions [8, 17, 22].
Most of the aforementioned works consider the asymptotic error distribution of explicit methods for SDEs. To the best of our knowledge, in the existing literature, there are no results on the asymptotic error distribution of diffusion-implicit numerical methods (the implicitness is introduced via the diffusion term g) for (1.1), despite their superior stability and frequent applications in constructing structure-preserving algorithms. The aim of this paper is to fill this gap. More precisely, we study the asymptotic error distribution of a class of SRK methods of strong order 111 applied to (1.1). Generally speaking, when establishing the asymptotic error distribution of a numerical method for (1.1), one usually finds a continuous version of the numerical method and properly decomposes the normalized error into some dominated terms and remainder terms. Then Jacod’s theory on convergence in distribution of conditional Gaussian martingales (cf. [4, 9, 10]) will play a key role in studying the convergence in distribution of stochastic integrals in dominated terms. However, for a diffusion-implicit or fully implicit method, the implicitness introduced in the diffusion term makes it not easy to find a proper continuous version of the numerical method, since the adaptedness of integrands in stochastic integrals must be taken into account. In addition, diffusion-implicit methods usually introduce the truncated Brownian increments to ensure the solvability, which also makes some trouble for the continuization of numerical methods and dealing with the convergence in distribution of stochastic integrals.
1.2. Framework to derive the asymptotic error distribution
Aiming at presenting our arguments clearly for dealing with the asymptotic error distribution of diffusion-implicit numerical methods, we will focus on the scalar noise case (i.e., $m=1$) and multidimensional additive noise, and study the SRK methods (3.9) and (5.2). We note that our framework remains applicable to multidimensional multiplicative noise case, though at the cost of increased computational effort. In order to circumvent the difficulty of directly making the continuization of the SRK method (3.9), our strategy is to find a fully explicit numerical method which shares the same asymptotic error distribution as the SRK method. Such an explicit numerical method needs to approximate the exact solution of (1.1) with strong order 111 and approximate the numerical solution of the SRK method with strong order strictly greater than 111. We call this explicit numerical method an appurtenant method of the SRK method (3.9). In order to determine the appurtenant method, we generalize the fundamental strong convergence theorem [21, Theorem 2.1] to give the strong error order between any two Markov chains (see Theorem 2.1). Based on Theorem 2.1 and the deterministic Taylor expansion, the appurtenant method can be constructed. Thus, we indeed give a framework to study the asymptotic error distribution of a diffusion-implicit or fully implicit numerical method $\{Y_{n}\}_{n=0}^{N}$, which is summarized as follows.
Table 1. Framework 111 for establishing the asymptotic error distribution of $Y_{N}$.
Step 1. Determine the strong convergence order $p$ of $\{Y_{n}\}_{n=0}^{N}$.
Step 2. Find the explicit appurtenant method $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ based on Theorem 2.1
and Taylor expansion. Here, $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ is required to approximate the solution to
(1.1) with strong order $p$ and approximates $\{Y_{n}\}_{n=0}^{N}$ with strong order $q>p$.
Step 3. Construct a proper continuization of $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ and establish the limit
distribution of $N^{p}(\widetilde{Y}_{N}-Y(T))$, which is that of $N^{p}({Y}_{N}-Y(T))$.
In Framework 1, the asymptotic error distribution of ${Y}_{N}$ is the limit distribution of $N^{p}({Y}_{N}-Y(T))$, which equals to $N^{p}({Y}_{N}-\widetilde{Y}_{N})+N^{p}(\widetilde{Y}_{N}-Y(T))$. Note that $N^{p}({Y}_{N}-\widetilde{Y}_{N})$ converges to 0 in probability since ${Y}_{N}-\widetilde{Y}_{N}$ is of strong order $q>p$. Accordingly, $N^{p}({Y}_{N}-Y(T))$ and $N^{p}(\widetilde{Y}_{N}-Y(T))$ have the same limit distribution due to Slutzky’s theorem (cf. [11, Theorem 13.18]).
1.3. Main results and contributions
[…]
Let us summarize the main contributions of this work as follows.
  • We propose a practical and effective framework (Framework 1) for establishing the asymptotic error distribution of diffusion-implicit or fully-implicit numerical methods for SDEs. In particular, this framework is successfully applied to deriving the asymptotic error distribution of a class of fully implicit methods for SDEs, which, to the best of our knowledge, is reported here for the first time. It is worth mentioning that Framework 1 is promisingly applicable to higher-order methods for SDEs and implicit numerical methods for other types of stochastic systems.
  • We identify the key parameter reflecting the growth rate of the mean-square error of the SRK method by analyzing the properties of the limit distribution. Our results reveal that among the SRK methods of strong order 111, those of weak order 222 have the smallest mean-square errors after a long time. Especially, the asymptotic error distribution of the considered SRK methods takes the unified form, when endowed with second order weak convergence.
  • We are the first to uncover a novel relationship between the asymptotic error distribution of stochastic numerical methods and their weak convergence order. From the main results of the paper, we conjecture that among stochastic numerical methods with same strong order, those with higher weak convergence order will possess smaller strong error in the long-time simulations.
The rest of the article is organized as follows. We establish Theorem 2.1 for estimating the strong error order between two Markov chains in Section 2 and Theorem 2.3 about stable convergence for a class of SDEs. In addition, we establish the convergence in distribution for a class of SDEs in order to handle the asymptotic error distribution for the multiplicative noise case. Sections 3 and 4 are devoted to establishing the asymptotic error distribution of the SRK method for (1.1) with scalar noise, by using Framework 1. Section 5 gives the asymptotic error distribution of the SRK method for the case of additive noise. Several numerical experiments are performed to verify our theoretical analyses in Section 6. Finally, Section 7 gives the summary of this work.
2. Preliminaries
In this section, we focus on presenting two theorems on fundamental strong convergence results for numerical method and convergence in distribution for a special class of SDEs.
[…]
Denote by $\mathbf{C}(\mathbb{R}^{d})$ (resp. $\mathbf{C}^{k}(\mathbb{R}^{d})$) the space of continuous (resp. $k$th continuously differentiable) functions defined on $\mathbb{R}^{d}$. Let $\mathbf{C}_{b}^{k}(\mathbb{R}^{d})$, $k\in\mathbb{N}^{+}$, stand for the subset of $\mathbf{C}^{k}(\mathbb{R}^{d})$, consisting of functions whose derivatives up to order $k$ are bounded. For a real-valued function $f\in\mathbf{C}^{k}(\mathbb{R}^{d})$, denote by $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k})$ the $k$th order Gǎteaux derivative along the directions $\xi_{1},\ldots,\xi_{k}\,\in\mathbb{R}^{d}$. For an $\mathbb{R}^{m}$-valued function $f=(f_{1},\ldots,f_{m})^{\top}\in\mathbf{C}^{k}(\mathbb{R}^{d})$, denote $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k}):=(\mathcal{D}^{k}f_{1}(x)(\xi_{1},\ldots,\xi_{k}),\ldots,\mathcal{D}^{k}f_{m}(x)(\xi_{1},\ldots,\xi_{k}))^{\top}$. Let $\mathbf{F}$ be the set of functions growing at most polynomially, i.e., a (tensor-valued) function $f\in\mathbf{F}$ means that there exist $C>0$ and $\eta>0$ such that for any $x\in\mathbb{R}^{d}$, $|f(x)|\leq C(1+|x|^{\eta})$ or $\|f(x)\|_{\otimes}\leq C(1+|x|^{\eta})$, where $\otimes$ denotes the norm of a tensor. Throughout this paper, let $K(a_{1},a_{2},...,a_{m})$ be some generic constant dependent on the parameters $a_{1},a_{2},...,a_{m}$ but independent of the step-size $h$, which may vary for each appearance.
2.1. Fundamental strong convergence theorem
In this part, we present a theorem for analyzing the strong error order between two Markov chains, which will be used to construct the appurtenant method of (3.9). Let $\{t_{0}<t_{1}<\cdots<t_{N}=T\}$, $N\in\mathbb{N}^{+}$ be a uniform partition of $[t_{0},T]$, i.e., $t_{k}-t_{k-1}=\frac{T-t_{0}}{N}:=h$ for any $k=1,\ldots,N$. Introduce two one-step mappings $\bar{X}_{t,x}(t+h)$ and $\widetilde{X}_{t,x}(t+h)$ depending on $t,x,h$, which are defined respectively by
$\bar{X}_{t,x}(t+h)$ $=x+\bar{A}(t,x,h;\mathbf{W}(s)-\mathbf{W}(t),t\leq s\leq t+h),$
$\widetilde{X}_{t,x}(t+h)$ $=x+\widetilde{A}(t,x,h;\mathbf{W}(s)-\mathbf{W}(t),t\leq s\leq t+h),$
where $\bar{A}(t,x,h;y)$ and $\widetilde{A}(t,x,h;y)$ are two measurable functions w.r.t. $(t,x,h;y)\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{m}$. Then one can construct two Markov chains $(\bar{X}_{k},\mathcal{F}_{t_{k}})$ and $(\widetilde{X}_{k},\mathcal{F}_{t_{k}})$, $k=0,1,\ldots,N$: Let $\bar{X}_{0}=\widetilde{X}_{0}\in\mathbf{L}^{2}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$, and for $k=0,\ldots,N-1$:
$\bar{X}_{k+1}=\bar{X}_{t_{k},\bar{X}_{k}}(t_{k+1})=\bar{X}_{k}+\bar{A}(t_{k},\bar{X}_{k},h;\mathbf{W}(s)-\mathbf{W}(t_{k}),\;t_{k}\leq s\leq t_{k+1}),~k=0,\ldots,N-1,$
$\widetilde{X}_{k+1}=\widetilde{X}_{t_{k},\widetilde{X}_{k}}(t_{k+1})=\widetilde{X}_{k}+\widetilde{A}(t_{k},\widetilde{X}_{k},h;\mathbf{W}(s)-\mathbf{W}(t_{k}),\;t_{k}\leq s\leq t_{k+1}),~k=0,\ldots,N-1.$
Authors of [21] proposed a fundamental strong convergence theorem, [21, Theorem 2.1], for numerical methods applied to SDEs. Following the arguments of proving [21, Theorem 2.1], we give the strong error order between $\bar{X}_{k}$ and $\widetilde{X}_{k}$.
Theorem 2.1.
Let $\{\bar{X}_{k}\}_{k=0}^{N}$ and $\{\widetilde{X}_{k}\}_{k=0}^{N}$ be defined as above. Assume that the following conditions hold:
  • (A1)
    For some $p\geq 1$, there exist $p_{2}\geq\frac{1}{2}$, $p_{1}\geq p_{2}+\frac{1}{2}$, $\alpha\geq 1$, $h_{0}\in(0,1]$ and $K_{1}>0$ such that for all $t\in[t_{0},T-h]$, $x\in\mathbb{R}^{d}$ and $h\in(0,h_{0}]$,
    $\big{|}\mathbf{E}\big{(}\bar{X}_{t,x}(t+h)-\widetilde{X}_{t,x}(t+h)\big{)}\big{|}$ $\leq K_{1}(1+|x|^{2\alpha})^{1/2}h^{p_{1}},$
    $\big{[}\mathbf{E}\big{|}\bar{X}_{t,x}(t+h)-\widetilde{X}_{t,x}(t+h)\big{|}^{2p}\big{]}^{1/(2p)}$ $\leq K_{1}(1+|x|^{2\alpha p})^{1/(2p)}h^{p_{2}}.$
  • (A2)
    For any $q\geq 1$, there is $K_{2}(q)>0$ such that for all $h\in(0,h_{0}]$ and $k=0,1,\ldots,N$,
    $\mathbf{E}|\bar{X}_{k}|^{q}+\mathbf{E}|\widetilde{X}_{k}|^{q}\leq K_{2}(q).$
  • (A3)
    There is a representation
    $\widetilde{X}_{t,x}(t+h)-\widetilde{X}_{t,y}(t+h)=x-y+V$
    for which
    $\mathbf{E}\big{|}\widetilde{X}_{t,x}(t+h)-\widetilde{X}_{t,y}(t+h)\big{|}^{2p}\leq|x-y|^{2p}(1+Kh),$
    $\mathbf{E}|V|^{2p}\leq K(1+|x|^{p\beta}+|y|^{p\beta})|x-y|^{2p}h^{p},$
    where $\beta\geq 1$ is a constant.
Then, for any $h\in(0,h_{0}]$ and $k=0,\ldots,N$,
$\big{[}\mathbf{E}\big{|}\bar{X}_{k}-\widetilde{X}_{k}\big{|}^{2p}\big{]}^{1/(2p)}\leq Kh^{p_{2}-1/2}.$
Proof.
This proof follows the same steps as in the proof of [21, Theorem 2.1], by replacing $X$ therein by $\widetilde{X}$ here. We note that [21, Assumption 2.1] is used to establish [21, Lemma 2.2] and the moment boundedness of $X(t_{k})$ therein. In fact, in the proof of [21, Theorem 2.1], only the one-step approximation error [21, Eq. (2.9)-(2.10)], the moment uniform boundedness of $X(t_{k})$ and $\bar{X}_{k}$ therein, and [21, Lemma 2.2] are used. These three conditions correspond to (A1), (A2) and (A3), respectively. Thus, the proof of this theorem can be completed by following the proof procedure of [21, Theorem 2.1]. ∎
2.2. Stable convergence in distribution
In this part, we introduce the notion of stable convergence in distribution and giving a theorem on convergence in distribution of solutions to a sequence of SDEs perturbed by processes stably converging in law.
Let $\{X_{n}\}_{n=1}^{\infty}$ be a sequence of random variables with values in a Polish space $E$, all defined on $(\Omega,\mathcal{F},\mathbf{P})$. Let $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}})$ be an extension of $(\Omega,\mathcal{F},\mathbf{P})$ and let $X$ be an $E$-valued random variable on the extension. We say that $X_{n}$ stably converges in law to $X$ in $E$, written $X_{n}\overset{stably}{\Longrightarrow}X$ in $E$, if
$\mathbf{E}[Zf(X_{n})]\to\tilde{\mathbf{E}}[Zf(X)]$
[…]
Proposition 2.2.
[10, Lemma 2.1] Let $Y$ be a random variable on another Polish space $F$. If $X_{n}\overset{stably}{\Longrightarrow}X$ in $E$, then $(Y,X_{n})\overset{stably}{\Longrightarrow}(Y,X)$ in $F\times E$.
On basis of Proposition 2.2, we present the following convergence in distribution of a family of solutions to systems of SDEs.
Theorem 2.3.
Let $\Phi^{N}=\{\Phi^{N}(t),t\in[0,T]\}$, $N\geq 1$ be $\{\mathcal{F}_{t}\}$-adapted $\mathbb{R}^{d_{2}}$-valued processes defined on $(\Omega,\mathcal{F},\mathbf{P})$, and $Y=\{Y(t),t\in[0,T]\}$ be an $\{\mathcal{F}_{t}\}$-adapted $\mathbb{R}^{d_{1}}$-valued process defined on $(\Omega,\mathcal{F},\mathbf{P})$ with almost sure continuous trajectories. Let $Z^{N}\in\mathbf{C}([0,T];\mathbb{R}^{d_{2}})$, $N\geq 1$, be the strong solution of
$Z^{N}(t)=\int_{0}^{t}B_{0}(Y(s))Z^{N}(s)\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(s))Z^{N}(s)\mathrm{d}W_{i}(s)+\Phi^{N}(t),~t\in[0,T],$ (2.1)
where $B_{i}:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}\times d_{2}}$ are bounded and $\mathcal{D}B_{i}\in\mathbf{F}$ (i.e., $\mathcal{D}B_{i}$ grow at most polynomially) for $i=0,\ldots,m$. Assume that the following conditions hold.
  • (1)
    There is $\beta>0$ such that $Y\in\mathbf{C}([0,T];\mathbf{L}^{p}(\Omega;\mathbb{R}^{d_{1}}))$ and $\|Y(t)-Y(s)\|_{\mathbf{L}^{p}(\Omega)}\leq K|t-s|^{\beta}$ for sufficiently large $p$.
  • (2)
    $\Phi^{N}\overset{stably}{\Longrightarrow}\Phi$ in $\mathbf{C}([0,T];\mathbb{R}^{d_{2}})$, where $\Phi$ is an $\{\tilde{\mathcal{F}_{t}}\}$-adapted process defined on $(\tilde{\Omega},\tilde{\mathcal{F}},\{\tilde{\mathcal{F}_{t}}\}_{t\in[0,T]},\tilde{\mathbf{P}})$, an extension of $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in[0,T]},\mathbf{P})$.
  • (3)
    $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|\Phi^{N}(t)|^{4}+\sup\limits_{t\in[0,T]}\tilde{\mathbf{E}}|\Phi(t)|^{4}<\infty$, and there is $\alpha>0$ such that $\sup\limits_{N\geq 1}\mathbf{E}|\Phi^{N}(t)-\Phi^{N}(s)|^{2}+\tilde{\mathbf{E}}|\Phi(t)-\Phi(s)|^{2}\leq K|t-s|^{\alpha}$ for any $t,s\in[0,T]$.
[…]
$Z(t)=\int_{0}^{t}B_{0}(Y(s))Z(s)\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(s))Z(s)\mathrm{d}W_{i}(s)+\Phi(t),~t\in[0,T].$ (2.2)
Proof.
The main trouble lies in that $Z^{N}$ is not a continuous function w.r.t. $(Y,W_{1},\ldots,W_{m},\Phi^{N})$. Thus, we will prove the conclusion based on the approximation argument for convergence in distribution; see [6, Theorem 3.2].
It follows from the boundedness of $B_{i}$, $i=0,1,\ldots,m$, and $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|\Phi^{N}(t)|^{4}<\infty$ that $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|Z^{N}(t)|^{4}<\infty$, which together with $\sup\limits_{N\geq 1}\mathbf{E}|\Phi^{N}(t)-\Phi^{N}(s)|^{2}\leq K|t-s|^{\alpha}$ yields
$\sup_{N\geq 1}\mathbf{E}|Z^{N}(t)-Z^{N}(s)|^{2}\leq K|t-s|^{\min(1,\alpha)}.$ (2.3)
Let $Z^{N,n}=\{Z^{N,n}(t),t\in[0,T]\}$ be the strong solution of
$Z^{N,n}(t)=\int_{0}^{t}B_{0}(Y(\kappa_{n}(s)))Z^{N,n}(\kappa_{n}(s))\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(\kappa_{n}(s)))Z^{N,n}(\kappa_{n}(s))\mathrm{d}W_{i}(s)+\Phi^{N}(t),$
[…]
$\;\mathbf{E}\|Z^{N,n}-Z^{N}\|_{\mathbf{C}([0,t])}^{2}$
$\leq$ $\;K\mathbf{E}\int_{0}^{t}|Z^{N,n}(\kappa_{n}(s))-Z^{N}(s)|^{2}\mathrm{d}s+K\sum_{i=0}^{m}\mathbf{E}\int_{0}^{t}|B_{i}(Y(\kappa_{n}(s)))-B_{i}(Y(s))|^{2}|Z^{N}(s)|^{2}\mathrm{d}s$
$\leq$ $\;K\mathbf{E}\int_{0}^{t}|Z^{N,n}(\kappa_{n}(s))-Z^{N}(\kappa_{n}(s))|^{2}\mathrm{d}s+K\mathbf{E}\int_{0}^{t}|Z^{N}(\kappa_{n}(s))-Z^{N}(s)|^{2}\mathrm{d}s$
$\;+K\sum_{i=0}^{m}\int_{0}^{t}\|B_{i}(Y(\kappa_{n}(s)))-B_{i}(Y(s))\|_{\mathbf{L}^{4}(\Omega)}^{2}\|Z^{N}(s)\|_{\mathbf{L}^{4}(\Omega)}^{2}\mathrm{d}s.$ (2.4)
Since $\mathcal{D}B_{i}\in\mathbf{F}$, $i=0,1,\ldots,m$, we have $|B_{i}(y_{1})-B_{i}(y_{2})|\leq K(1+|y_{1}|^{l}+|y_{2}|^{l})|y_{1}-y_{2}|$ for some $l>0$. Thus, the Hölder inequality and the assumption on $Y$ gives
$\sum_{i=0}^{m}\|B_{i}(Y(\kappa_{n}(s)))-B_{i}(Y(s))\|_{\mathbf{L}^{4}(\Omega)}^{2}\leq Kn^{-2\beta}.$
Plugging the above relation into (2.2) and using $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|Z^{N}(t)|^{4}<\infty$ and (2.3), we arrive at
$\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,t])}\leq K\int_{0}^{t}\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,s])}\mathrm{d}s+Kn^{-\min(1,\alpha,2\beta)}$
with $K$ being independent of $N$. In this way, one has $\lim\limits_{n\to\infty}\sup\limits_{N\geq 1}\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,T])}=0$ due to the Gronwall inequality. This indicates $Z^{N,n}$ and $Z^{N}$ satisfies the condition (A1) of [6, Theorem 3.2]
For any fixed $n\in\mathbb{N}^{+}$, define the mapping $\Gamma^{n}:\mathbf{C}([0,T];\mathbb{R}^{d_{1}})\times\mathbf{C}([0,T];\mathbb{R})^{\otimes m}\times\mathbf{C}([0,T];\mathbb{R}^{d_{2}})\to\mathbf{C}([0,T];\mathbb{R}^{d_{2}})$ which maps $(u,g_{1},\ldots,g_{m},h)$ to the solution of
$f(t)=\int_{0}^{t}B_{0}(u(\kappa_{n}(s)))f(\kappa_{n}(s))\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(u(\kappa_{n}(s)))f(\kappa_{n}(s))\mathrm{d}g_{i}(s)+h(t),~t\in[0,T].$
[…]
$Z^{\infty,n}(t)=\int_{0}^{t}B_{i}(Y(\kappa_{n}(s)))Z^{\infty,n}(\kappa_{n}(s))\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(\kappa_{n}(s)))Z^{\infty,n}(\kappa_{n}(s))\mathrm{d}W_{i}(s)+\Phi(t),~t\in[0,T].$
It follows from the assumption on $\Phi$ that $\sup\limits_{t\in[0,T]}\tilde{\mathbf{E}}|Z(t)|^{4}<\infty$ and $\tilde{\mathbf{E}}|Z(t)-Z(s)|^{2}\leq K|t-s|^{1\wedge\alpha}$. Similarly to the estimate for $\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,T])}$, one can show that $\tilde{\mathbf{E}}\|Z^{\infty,n}-Z\|^{2}_{\mathbf{C}([0,T])}\leq Kn^{-\min(1,\alpha,2\beta)}$. In this way, we have
[…]
Thus, the conditions (A2) and (A3) of [6, Theorem 3.2] are fulfilled. Thus, the proof is complete by applying [6, Theorem 3.2]. ∎
3. Properties of stochastic Runge–Kutta method for multiplicative noise
In this section, for the scalar multiplicative noise, we establish the solvability and strong convergence order of the SRK method, and construct the appurtenant method which shares the same asymptotic error distribution as the SRK method.
Consider the following SDE with scalar noise:
$\begin{cases}\mathrm{d}Y(t)=f(Y(t))\mathrm{d}t+g(Y(t))\circ\mathrm{d}W(t),~t\in(0,T],\\ Y(0)=Y_{0}\in\mathbb{R}^{d},\end{cases}$ (3.1)
where $f,g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ are Lipschitz continuous, $W$ is a one-dimensional standard Brownian motion, and $g(Y(t))\circ\mathrm{d}W(t)$ denotes the Stratonovich stochastic differential. When $g\in\mathbf{C}^{1}(\mathbb{R}^{d})$, (3.1) is equivalent to the following Itô-type SDE:
$\begin{cases}\mathrm{d}Y(t)=\bar{f}(Y(t))\mathrm{d}t+g(Y(t))\mathrm{d}W(t),~t\in(0,T],\\ Y(0)=Y_{0}\in\mathbb{R}^{d},\end{cases}$ (3.2)
where $\bar{f}(y)=f(y)+\frac{1}{2}\nabla g(y)g(y)$, $y\in\mathbb{R}^{d}$, and $g(Y(t))\mathrm{d}W(t)$ denotes the Itô stochastic differential. For convenience, we always assume that $Y_{0}$ is nonrandom.
Assumption 1.
Assume that $f$, $g$ and $(\nabla g)g$ are Lipschitz continuous.
It is well-known that under Assumption 1, (3.1) admits a unique strong solution which satisfies that for any $p\geq 2$, and $t,s\in[0,T]$,
$\|Y(t)\|_{\mathbf{L}^{p}(\Omega)}$ $\leq C_{1}(p)e^{C_{1}(p)T}(1+|Y_{0}|),$ (3.3)
$\|Y(t)-Y(s)\|_{\mathbf{L}^{p}(\Omega)}$ $\leq K(p,T)|t-s|^{1/2},$ (3.4)
where $C_{1}(p)$ depends on $p$ but independent of $T$ (cf. [15, Chapter 2.4]).
3.1. Stochastic Runge–Kutta method
Generally speaking, in order to ensure the solvability and stability of diffusion-implicit methods, one needs to use the truncated random variables $\Delta\widehat{W}_{n}$ instead of $\Delta W_{n}=W(t_{n+1})-W(t_{n})$, $n=0,\ldots,N-1$. In detail, we represent $\Delta W_{n}=\sqrt{h}\xi_{n}$, where $\xi_{0},\xi_{1},\ldots,\xi_{N-1}$ are independent $\mathcal{N}(0,1)$-distributed random variables. Then define $\Delta\widehat{W}_{n}:=\sqrt{h}\zeta_{n,h}$ by
$\zeta_{n,h}:=\begin{cases}\xi_{n},\quad&|\xi_{n}|\leq A_{h},\\ A_{h},\quad&\xi_{n}>A_{h},\\ -A_{h},\quad&\xi_{n}<-A_{h},\end{cases}$ (3.5)
where $A_{h}:=\sqrt{2\kappa|\ln h|}$ with $\kappa\geq 1$ being a constant. Then one has the following properties:
$\mathbf{E}|\xi_{n}-\zeta_{n,h}|^{p}$ $\leq Kh^{\kappa},~\forall~p\geq 1,$ (3.6)
$\|\xi_{n}^{m}-\zeta_{n,h}^{m}\|_{\mathbf{L}^{p}(\Omega)}$ $\leq K(p,m,\varepsilon)h^{\kappa/p-\varepsilon},~\forall~m\in\mathbb{N}^{+},~p\geq 1,~\varepsilon\in(0,\frac{\kappa}{p}),$ (3.7)
$|\mathbf{E}(\xi_{n}^{m}-\zeta_{n,h}^{m})|$ $\leq K(m,\varepsilon)h^{\kappa-\varepsilon},~\forall~m\in\mathbb{N}^{+},~\varepsilon\in(0,\kappa).$ (3.8)
In fact, (3.6) can be proved similarly to [16, Page 39, Eq. (3.36)]. Note that $\mathbf{E}|\zeta_{n,h}|^{q}\leq\mathbf{E}|\xi_{n}|^{q}\leq K(q)$ for any $q\geq 1$. By the Hölder inequality, for any $\varepsilon\in(0,\frac{\kappa}{p})$, we have that $\mathbf{E}|\xi_{n}^{m}-\zeta_{n,h}^{m}|^{p}=\mathbf{E}\big{(}|m\int_{0}^{1}(\zeta_{n,h}+\lambda(\xi_{n}-\zeta_{n,h}))^{m-1}\mathrm{d}\lambda|^{p}|\xi_{n}-\zeta_{n,h}|^{p}\big{)}\leq\big{(}\mathbf{E}|\xi_{n}-\zeta_{n,h}|^{\frac{p}{1-\varepsilon p/\kappa}}\big{)}^{1-\varepsilon p/\kappa}K(p,\varepsilon,m)\leq K(p,\varepsilon,m)h^{\kappa-p\varepsilon}$, which verifies (3.7). In addition, (3.8) is a direct result of (3.7) with $p=1$.
Consider the following $s_{0}$-stage SRK for (3.1)
$\begin{cases}Z_{n,i}=Y_{n}^{SRK}+h\sum\limits_{j=1}^{s_{0}}a_{ij}f(Z_{n,j})+\Delta\widehat{W}_{n}\sum\limits_{j=1}^{s_{0}}b_{ij}g(Z_{n,j}),~i=1,\ldots,s_{0},\\ Y^{SRK}_{n+1}=Y^{SRK}_{n}+h\sum\limits_{i=1}^{s_{0}}\alpha_{i}f(Z_{n,i})+\Delta\widehat{W}_{n}\sum\limits_{i=1}^{s_{0}}\beta_{i}g(Z_{n,i}),~n=0,\ldots,N-1,\end{cases}$ (3.9)
with $Y_{0}^{SRK}=Y_{0}$. The coefficients of (3.9) can be represented by the Butcher tableau of the form   $A$ $B$ $\alpha^{\top}$ $\beta^{\top}$ , where $A=(a_{ij})$, $B=(b_{ij})$, $\alpha=(\alpha_{1},\ldots,\alpha_{s_{0}})^{\top}$ and $\beta=(\beta_{1},\ldots,\beta_{s_{0}})^{\top}$. Then the one-step approximation of (3.9) can be written as
$Y^{SRK}_{t,y}(t+h):=y+h\sum\limits_{i=1}^{s_{0}}\alpha_{i}f(Z_{i})+\Delta\widehat{W}\sum\limits_{i=1}^{s_{0}}\beta_{i}g(Z_{i}).$ (3.10)
Here $\Delta\widehat{W}=\sqrt{h}\zeta$ is the truncation of $W(t+h)-W(t)=\sqrt{h}\xi$ for which $\xi\sim\mathcal{N}(0,1)$ and $\zeta$ is defined as in (3.5) with $\xi_{n}$ replaced by $\xi$. In addition, $Z=(Z_{1}^{\top},\ldots,Z_{s_{0}}^{\top})^{\top}$ is determined by
$Z=(e\otimes I_{d})y+h(A\otimes I_{d})F(Z)+\Delta\widehat{W}(B\otimes I_{d})G(Z),$ (3.11)
where $F(Z)=(f(Z_{1})^{\top},\ldots,f(Z_{s_{0}})^{\top})^{\top}$, $G(Z)=(g(Z_{1})^{\top},\ldots,g(Z_{s_{0}})^{\top})^{\top}$, every component of $e\in\mathbb{R}^{s_{0}}$ equals to 111, $I_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\otimes$ denotes the Kronecker product of matrices. The following lemma gives the solvability of (3.11).
Lemma 3.1.
Let Assumption 1 hold. Then there is $h_{1}>0$ such that for any $h\in(0,h_{1}]$, $y\in\mathbb{R}^{d}$, $\omega\in\Omega$, (3.11) is uniquely solvable. Moreover, there is $C_{0}>0$ such that
$|Z_{i}-y|\leq C_{0}(1+|y|)(h+|\Delta\widehat{W}|),~h\in(0,h_{1}].$
Since $h_{1}$ in Lemma 3.1 is independent of $y$ and $\omega$, we obtain the well-posedness of (3.9). Next, we present the moment boundedness of (3.9).
Lemma 3.2.
Let Assumption 1 hold. Then for any $h\leq h_{1}\wedge 1$, we have that for any $p\geq 1$,
$\sup_{n=0,1,\ldots,N}\mathbf{E}|Y^{SRK}_{n}|^{p}\leq K(p,T).$
We postpone the proof of Lemmas 3.1 and 3.2 to the appendix. In the subsequent text of this section, we always assume that $h\leq h_{1}\wedge 1$ without extra statement. The first order mean-square convergence condition of (3.9) have been obtained in [3, Eq. (3.7)]. Here, we generalize it to the $\|\cdot\|_{L^{2p}(\Omega)}$-norm for $p\in[1,\frac{\kappa}{2}]$.
Theorem 3.3.
Let Assumption 1 hold. Let $f\in\mathbf{C}^{2}(\mathbb{R}^{d})$, $g\in\mathbf{C}^{3}(\mathbb{R}^{d})$ and $\mathcal{D}^{2}f,\,\mathcal{D}^{3}g\in\mathbf{F}$. If $\alpha^{\top}e=\beta^{\top}e=1$, $\beta^{\top}(Be)=\frac{1}{2}$ and $\kappa\geq 2$, then for any $p\in[1,\frac{\kappa}{2}]$,
$\sup_{n=0,1,\ldots,N}\|Y^{SRK}_{n}-Y(t_{n})\|_{\mathbf{L}^{2p}(\Omega)}\leq Kh.$
Theorem 3.3 can be proved based on Theorem 2.1 and (3.6)-(3.8), whose proof is postponed to the appendix.
3.2. An appurtenant numerical method
For $a=(a_{1},\ldots,a_{d})^{\top},~b=(b_{1},\ldots,b_{d})^{\top}\in\mathbb{R}^{d}$, denote $a*b\in\mathbb{R}^{d}$ with $(a*b)_{i}=a_{i}b_{i}$, $i=1,\ldots,d$ and $a^{m}\in\mathbb{R}^{d}$ with $(a^{m})_{i}=a_{i}^{m}$. The following lemma gives a sufficient expansion of $Y_{t,y}^{SRK}(t+h)$, helping us to find the appurtenant numerical method as in Framework 1, whose proof is deferred to the appendix.
Lemma 3.4.
Let Assumption 1 hold. Let $f\in\mathbf{C}^{3}(\mathbb{R}^{d})$, $g\in\mathbf{C}^{4}(\mathbb{R}^{d})$ and $\mathcal{D}^{3}f,\,\mathcal{D}^{4}g\in\mathbf{F}$. Moreover, let $\alpha^{\top}e=\beta^{\top}e=1$, $\beta^{\top}(Be)=\frac{1}{2}$. Then $Y^{SRK}_{t,y}(t+h)$ has following representation
$Y^{SRK}_{t,y}(t+h)=$ $\;y+\Delta\widehat{W}g(y)+hf(y)+\frac{1}{2}\Delta\widehat{W}^{2}\nabla g(y)g(y)+\Delta\widehat{W}hF_{1}(y)+\Delta\widehat{W}^{3}F_{2}(y)$
$\;+h^{2}\alpha^{\top}(Ae)\nabla f(y)f(y)+\Delta\widehat{W}^{2}hF_{3}(y)+\Delta\widehat{W}^{4}F_{4}(y)+R^{2}_{Y^{SRK}},$ (3.12)
where
$F_{1}(y)=$ $\;\alpha^{\top}(Be)\nabla f(y)g(y)+\beta^{\top}(Ae)\nabla g(y)f(y),$
$F_{2}(y)=$ $\;\beta^{\top}(B(Be))(\nabla g(y))^{2}g(y)+\frac{1}{2}\beta^{\top}(Be)^{2}\mathcal{D}^{2}g(y)(g(y),g(y)),$
$F_{3}(y)=$ $\;\alpha^{\top}(B(Be))\nabla f(y)\nabla g(y)g(y)+\beta^{\top}(A(Be))\nabla g(y)\nabla f(y)g(y)+\beta^{\top}(B(Ae))(\nabla g(y))^{2}f(y)$
$\;+\frac{1}{2}\alpha^{\top}(Be)^{2}\mathcal{D}^{2}f(y)(g(y),g(y))+\beta^{\top}((Ae)*(Be))\mathcal{D}^{2}g(y)(f(y),g(y)),$
$F_{4}(y)=$ $\;\beta^{\top}(B(B(Be)))(\nabla g(y))^{3}g(y)+\frac{1}{2}\beta^{\top}(B(Be)^{2})\nabla g(y)\mathcal{D}^{2}g(y)(g(y),g(y))$
$\;+\beta^{\top}((Be)*(B(Be)))\mathcal{D}^{2}g(y)(g(y),\nabla g(y)g(y))+\frac{1}{6}\beta^{\top}(Be)^{3}\mathcal{D}^{3}g(y)(g(y),g(y),g(y)),$
and there is $\alpha^{\prime}\geq 1$ such that for any $p\geq 1$, $\|R^{2}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\alpha^{\prime}})h^{5/2}$.
According to (3.12), we can define the candidate of the appurtenant method of the SRK method (3.9). For this end, let $\bar{Y}_{t,y}(t+h)$ be the one-step approximation defined by
$\bar{Y}_{t,y}(t+h)=$ $\;y+\Delta Wg(y)+hf(y)+\frac{1}{2}\Delta W^{2}\nabla g(y)g(y)+\Delta WhF_{1}(y)+\Delta W^{3}F_{2}(y)$
$\;+h^{2}\alpha^{\top}(Ae)\nabla f(y)f(y)+\Delta W^{2}hF_{3}(y)+\Delta W^{4}F_{4}(y).$ (3.13)
Then, we have following properties.
Lemma 3.5.
Let the assumptions of Lemma 3.4 hold. If $\kappa\geq 3$, then there is some $\alpha^{\prime\prime}\geq 1$ such that for $0\leq t\leq T-h$ and $y\in\mathbb{R}^{d}$,
$\big{|}\mathbf{E}(Y^{SRK}_{t,y}(t+h)-\bar{Y}_{t,y}(t+h))\big{|}$ $\leq K(1+|y|^{\alpha^{\prime\prime}})h^{5/2},$
$(\mathbf{E}|Y^{SRK}_{t,y}(t+h)-\bar{Y}_{t,y}(t+h)|^{2})^{1/2}$ $\leq K(1+|y|^{\alpha^{\prime\prime}})h^{2}.$
Proof.
By (3.12) and (3.2), one has
$\;Y^{SRK}_{t,y}(t+h)-\bar{Y}_{t,y}(t+h)=(\Delta\widehat{W}-\Delta W)g(y)+\frac{1}{2}(\Delta\widehat{W}^{2}-\Delta W^{2})\nabla g(y)g(y)+(\Delta\widehat{W}-\Delta W)hF_{1}(y)$
$\;+(\Delta\widehat{W}^{3}-\Delta W^{3})F_{2}(y)+(\Delta\widehat{W}^{2}-\Delta W^{2})hF_{3}(y)+(\Delta\widehat{W}^{4}-\Delta W^{4})F_{4}(y)+R^{2}_{Y^{SRK}}.$
Using (3.8) gives $|\mathbf{E}(\Delta\widehat{W}^{m}-\Delta W^{m})|\leq Kh^{\frac{m}{2}+\kappa-\varepsilon}\leq Kh^{4-\varepsilon}$ for $m\geq 2$ and $\kappa\geq 3$. Applying (3.6), we obtain $(\mathbf{E}(\Delta\widehat{W}-\Delta W)^{2})^{1/2}\leq Kh^{\frac{\kappa+1}{2}}\leq Kh^{2}$. Further, it follows from (3.7) that $(\mathbf{E}(\Delta\widehat{W}^{m}-\Delta W^{m})^{2})^{1/2}\leq Kh^{\frac{m+\kappa}{2}-\varepsilon}\leq Kh^{5/2-\varepsilon}$ for $m\geq 2$. Combining the above estimates, $\|R^{2}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\alpha^{\prime}})h^{5/2}$ (see Lemma 3.4), and $F_{i}\in\mathbf{F}$ for $i=1,2,3,4$, we complete the proof. ∎
Let $\{\bar{Y}_{n}\}_{n=0}^{N}$ be the numerical solution generated by $\bar{Y}_{t,h}(t+h)$, i.e., $\bar{Y}_{n+1}=\bar{Y}_{t_{n},\bar{Y}_{n}}(t_{n+1})$, $n=0,\ldots,N-1$ and $\bar{Y}_{0}=Y_{0}$. Based on Lemma 3.5 and Theorem 2.1, it can be shown that the mean-square error order between $\bar{Y}_{n}-Y^{SRK}_{n}$ is 1.51.51.5, and $\{\bar{Y}_{n}\}_{n=0}^{N}$ has mean-square order 111 when approximating (3.1). This is to say, $\{\bar{Y}_{n}\}_{n=0}^{N}$ is a candidate of the appurtenant method of the SRK method (3.9). However, the terms $\Delta W^{i}$, $i=2,3,4$ in (3.2) are very inconvenient for us to give a proper continuous version of $\{\bar{Y}_{n}\}_{n=0}^{N}$ for studying its asymptotic error distribution. For this end, we use the multiple Itô integral to represent $\Delta W^{i}$, $i=2,3,4$, and give the appurtenant method of (3.9).
First, it holds that
$\Delta W^{2}=(W(t+h)-W(t))^{2}=2\int_{t}^{t+h}(W(r)-W(t))\mathrm{d}W(r)+h.$ (3.14)
Then by the Itô formula,
$(W(s)-W(t))^{3}=3\int_{t}^{s}(W(r)-W(t))^{2}\mathrm{d}W(r)+3\int_{t}^{s}(W(r)-W(t))\mathrm{d}r,$
which, along with (3.14), yields
$(W(s)-W(t))^{3}=6\int_{t}^{s}\int_{t}^{r}(W(u)-W(t))\mathrm{d}W(u)\mathrm{d}W(r)+3(s-t)(W(s)-W(t)).$
Accordingly,
$\Delta W^{3}=6\int_{t}^{t+h}\int_{t}^{r}(W(u)-W(t))\mathrm{d}W(u)\mathrm{d}W(r)+3h\Delta W.$ (3.15)
Similarly, applying the Itô formula and (3.14), one has
$\Delta W^{4}=4\int_{t}^{t+h}(W(s)-W(t))^{3}\mathrm{d}W(s)+12\int_{t}^{t+h}\int_{t}^{s}(W(r)-W(t))\mathrm{d}W(r)\mathrm{d}s+3h^{2}.$ (3.16)
Plugging (3.14)-(3.16) into (3.2) leads to
$\;\bar{Y}_{t,y}(t+h)=y+\Delta Wg(y)+h\bar{f}(y)+\nabla g(y)g(y)\int_{t}^{t+h}(W(s)-W(t))\mathrm{d}W(s)$
$\;+\Delta Wh(F_{1}(y)+3F_{2}(y))+6F_{2}(y)\int_{t}^{t+h}\int_{t}^{s}(W(r)-W(t))\mathrm{d}W(r)\mathrm{d}W(s)$
$\;+h^{2}[\alpha^{\top}(Ae)\nabla f(y)f(y)+F_{3}(y)+3F_{4}(y)]+R_{\bar{Y}},$ (3.17)
where $\bar{f}$ has been defined in (3.2) and
$R_{\bar{Y}}:=$  $\;2h\int_{t}^{t+h}(W(s)-W(t))\mathrm{d}W(s)F_{3}(y)+F_{4}(y)\Big{[}4\int_{t}^{t+h}(W(s)-W(t))^{3}\mathrm{d}W(s)$
$\;+12\int_{t}^{t+h}\int_{t}^{s}(W(r)-W(t))\mathrm{d}W(r)\mathrm{d}s\Big{]}.$
Ignoring the remainder $R_{\bar{Y}}$, we introduce the one-step approximation $\widetilde{Y}_{t,y}(t+h)$:
$\;\widetilde{Y}_{t,y}(t+h)=y+\Delta Wg(y)+h\bar{f}(y)+\nabla g(y)g(y)\int_{t}^{t+h}(W(s)-W(t))\mathrm{d}W(s)$
$\;+\Delta Wh(F_{1}(y)+3F_{2}(y))+6F_{2}(y)\int_{t}^{t+h}\int_{t}^{s}(W(r)-W(t))\mathrm{d}W(r)\mathrm{d}W(s)$
$\;+h^{2}[\alpha^{\top}(Ae)\nabla f(y)f(y)+F_{3}(y)+3F_{4}(y)].$ (3.18)
Further, define the numerical solution $\{\widetilde{Y}_{n}\}_{n=0}^{N}$:
$\widetilde{Y}_{n+1}=\widetilde{Y}_{t_{n},\widetilde{Y}_{n}}(t_{n+1}),~n=0,\ldots,N-1,~\widetilde{Y}_{0}=Y_{0}.$ (3.19)
Next we show that (3.19) is an appurtenant method of the SRK method (3.9). Denote =$\mathcal{H}=\{f,g,(\nabla f)f,(\nabla f)g,(\nabla g)f,(\nabla g)g,(\nabla g)^{2}f,(\nabla g)^{2}g,(\nabla f\nabla g)g,(\nabla g\nabla f)g,\mathcal{D}^{2}f(g,g),\mathcal{D}^{2}g(g,g),\\ \mathcal{D}^{2}g(f,g),(\nabla g)^{3}g,\nabla g\big{(}\mathcal{D}^{2}g(g,g)\big{)},\mathcal{D}^{2}g(g,(\nabla g)g),\mathcal{D}^{3}g(g,g,g)\}$.
Lemma 3.6.
Let the assumptions of Lemma 3.4 hold and $\kappa\geq 3$. If every function from $\mathcal{H}$ is Lipschitz continuous, then
$\sup_{n=0,\ldots,N}(\mathbf{E}|Y^{SRK}_{n}-\widetilde{Y}_{n}|^{2})^{1/2}\leq Kh^{3/2}.$
Proof.
By (3.17) and (3.18), one has
$\mathbf{E}(\bar{Y}_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h))=0,~(\mathbf{E}|\bar{Y}_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h)|^{2})^{1/2}\leq K(1+|y|)h^{2},$
which, together with Lemma 3.5 and the Minkowski inequality, yields
$\big{|}\mathbf{E}(Y^{SRK}_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h))\big{|}$ $\leq K(1+|y|^{\alpha^{\prime\prime}})h^{5/2},$
$(\mathbf{E}|Y^{SRK}_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h)|^{2})^{1/2}$ $\leq K(1+|y|^{\alpha^{\prime\prime}})h^{2}.$
This verifies the condition (A1) of Theorem 2.1.
Denote $P(y)=\widetilde{Y}_{t,y}(t+h)-y-\Delta Wg(y)$. Under the assumptions on $f$ and g, we have that $\bar{f}$, $\nabla gg$, $\nabla ff$ and $F_{i}$, $i=1,2,3,4$ are Lipschitz continuous and thus grow at most linearly. Hence, it follows from (3.18) that $|P(y)|\leq(1+|y|)M$ where $M$ is a non-negative random variable with $\|M\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h$ for $p\geq 1$. Then, one has
$|\mathbf{E}(\widetilde{Y}_{t,y}(t+h)-y)|=|\mathbf{E}(\Delta Wg(y))+\mathbf{E}P(y)|\leq K(1+|y|)h.$
In addition, $|\widetilde{Y}_{t,y}(t+h)-y|\leq K(1+|y|)|\Delta W|+(1+|y|)M\leq KM^{\prime}(1+|y|)h^{1/2}$, where $M^{\prime}:=h^{-1/2}(|\Delta W|+M)$. Since $\mathbf{E}|M^{\prime}|^{p}<\infty$ for any $p\geq 1$, we obtain that for any $q\geq 1$,
$\sup_{n=0,1,\ldots,N}\mathbf{E}|\widetilde{Y}_{n}|^{q}\leq K(q)$ (3.20)
due to [16, Page 102, Lemma 2.2]. This combined with Lemma 3.2 verifies the condition (A2) of Theorem 2.1.
It follows from the definition of $P$ and (3.18) that $|P(x)-P(y)|\leq K|x-y|M^{\prime\prime}$ with $\|M^{\prime\prime}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h$ for $p\geq 1$. Further, we have that $\widetilde{Y}_{t,x}(t+h)-\widetilde{Y}_{t,y}(t+h)=x-y+\Delta W(g(x)-g(y))+P(x)-P(y)$. Thus, a direct computation leads to
$\mathbf{E}|\widetilde{Y}_{t,x}(t+h)-\widetilde{Y}_{t,y}(t+h)|^{2}$ $\leq|x-y|^{2}(1+Kh).$
$\mathbf{E}|\Delta W(g(x)-g(y))+P(x)-P(y)|^{2}$ $\leq K|x-y|^{2}h.$
Thus the condition (A3) of Theorem 2.1 is also satisfied. Finally the proof is finished as a result of Theorem 2.1. ∎
Remark 3.7.
We remark that a sufficient condition, to make every function belonging to $\mathcal{H}$ Lipschitz continuous, is that $f$ and $g$ are bounded, $f\in\mathbf{C}_{b}^{3}(\mathbb{R}^{d})$, and $g\in\mathbf{C}_{b}^{4}(\mathbb{R}^{d})$. In addition, an example of unbounded $f$ and $g$ satisfying the conditions of Lemma 3.6 is $f(x)=g(x)=\ln(1+x^{2})$, $x\in\mathbb{R}$.
By Theorem 3.3 and Lemma 3.6, we have that $\widetilde{Y}_{n}$ converges to $Y(t_{n})$ of mean-square order 111. Thus, $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ is an appurtenant method of (3.9). In fact, we can prove that $\sup_{n=0,1,\ldots,N}\|\widetilde{Y}_{n}-Y(t_{n})\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h$ for any $p\geq 2$.
Lemma 3.8.
Let the assumptions of Lemma 3.4 hold. If every function from $\mathcal{H}$ is Lipschitz continuous, then for any $p\geq 1$,
$\sup_{n=0,1,\ldots,N}\|\widetilde{Y}_{n}-Y(t_{n})\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h.$
Proof.
It follows from (A.13) and (3.14) that
$Y_{t,y}(t+h)=y+\Delta Wg(y)+\bar{f}(y)h+\nabla g(y)g(y)\int_{t}^{t+h}(W(s)-W(t))\mathrm{d}W(s)+R^{1}_{Y}.$
Comparing $Y_{t,y}(t+h)$ and $\widetilde{Y}_{t,y}(t+h)$, one can prove $Y_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h)$ satisfies the condition (A1) of Theorem 2.1 for any $p\geq 1$ with $p_{1}=2$ and $p_{2}=\frac{3}{2}$. In addition, the condition (A2) of Theorem 2.1 holds due to (3.3) and (3.20). Also, $Y_{t,y}(t+h)$ fulfills the condition (A3) of Theorem 2.1; see [21, Lemma 2.2]. Thus, the proof is complete based on Theorem 2.1. ∎
For convenience of statement, we introduce the following assumption which implies Assumption 1.
Assumption 2.
Assume that $f\in\mathbf{C}^{3}(\mathbb{R}^{d})$, $g\in\mathbf{C}^{4}(\mathbb{R}^{d})$, $\mathcal{D}^{3}f,\,\mathcal{D}^{4}g\in\mathbf{F}$ and every function from $\mathcal{H}$ is Lipschitz continuous. Moreover, assume that $\alpha^{\top}e=\beta^{\top}e=1$, $\beta^{\top}(Be)=\frac{1}{2}$ and $\kappa\geq 3$.
Under the above assumption, $\widetilde{Y}_{N}$ and $Y^{SRK}_{N}$ have the same asymptotic error distribution.
Lemma 3.9.
Let Assumption 2 hold. Then $N(Y^{SRK}_{N}-Y(T))$ and $N(\widetilde{Y}_{N}-Y(T))$ have the same limit distribution as $N\to\infty$, if one of them converges in distribution.
Proof.
Note that $N(Y^{SRK}_{N}-Y(T))=N(\widetilde{Y}_{N}-Y(T))+Q_{N}$ with $Q_{N}=N(Y^{SRK}_{N}-\widetilde{Y}_{N})$. By Lemma 3.6, $Q_{N}$ converges to 0 in probability. Thus, the proof is finished as a result of Slutzky’s theorem. ∎
4. Asymptotic error distribution for multiplicative noise
4.1. Limit distribution of normalized error process
In this section, we give the asymptotic error distribution of $Y_{N}^{SRK}$, which is that of $\widetilde{Y}_{N}$ due to Lemma 3.9. For this end, we introduce the continuous version of $\{\widetilde{Y}_{n}\}_{n=0}^{N}$:
$\;\widetilde{Y}(t)=Y_{0}+\int_{0}^{t}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}s+\int_{0}^{t}g(\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}W(s)$
$\;+\int_{0}^{t}\int_{\kappa_{N}(s)}^{s}\nabla g(\widetilde{Y}(\kappa_{N}(r)))g(\widetilde{Y}(\kappa_{N}(r)))\mathrm{d}W(r)\mathrm{d}W(s)+h\int_{0}^{t}\big{[}F_{1}(\widetilde{Y}(\kappa_{N}(s)))+3F_{2}(\widetilde{Y}(\kappa_{N}(s)))\big{]}\mathrm{d}W(s)$
$\;+6\int_{0}^{t}\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}F_{2}(\widetilde{Y}(\kappa_{N}(u)))\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}W(s)+h\int_{0}^{t}\big{[}\alpha^{\top}(Ae)\nabla f(\widetilde{Y}(\kappa_{N}(s)))f(\widetilde{Y}(\kappa_{N}(s)))$
$\;+F_{3}(\widetilde{Y}(\kappa_{N}(s)))+3F_{4}(\widetilde{Y}(\kappa_{N}(s)))\big{]}\mathrm{d}s,$ (4.1)
where $\kappa_{N}(s)=\lfloor\frac{s}{h}\rfloor h=\lfloor\frac{Ns}{T}\rfloor\frac{T}{N}$. Then one has $\widetilde{Y}(t_{n})=\widetilde{Y}_{n}$, $n=0,\ldots,N$ from (3.19) and (4.1). It is not difficult to show that $\widetilde{Y}$ converges to $Y$ with strong order 111.
Lemma 4.1.
Let Assumption 2 hold. Then for any $p\geq 1$,
$\sup_{t\in[0,T]}\|Y(t)-\widetilde{Y}(t)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h.$
Proof.
By (4.1), we have that for any $t\in[t_{n},t_{n+1}]$,
$\widetilde{Y}(t)=\widetilde{Y}_{n}+\int_{t_{n}}^{t}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}s+\int_{t_{n}}^{t}g(\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}W(s)+P^{N}(t)$
with $\sup_{t\in[t_{n},t_{n+1}]}\|P^{N}(t)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h$ for any $p\geq 1$. Further, it holds that
$\widetilde{Y}(t)-Y(t)=$ $\;\widetilde{Y}_{n}-Y(t_{n})+\int_{t_{n}}^{t}\big{[}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))-\bar{f}(Y(s))\big{]}\mathrm{d}s$
$\;+\int_{t_{n}}^{t}\big{[}g(\widetilde{Y}(\kappa_{N}(s)))-g(Y(s))\big{]}\mathrm{d}W(s)+P^{N}(t),~t\in[t_{n},t_{n+1}].$
For any Lipschitz function $G$, we have $\|G(Y(s))-G(\widetilde{Y}(\kappa_{N}(s)))\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h^{1/2}$ for $p\geq 1$ due to (3.4) and Lemma 3.8. Thus, the proof is complete by the Minkowski inequality and BDG inequality. ∎
We are in the position to give an expansion for the normalized error $N(\widetilde{Y}(t)-Y(t))$.
Lemma 4.2.
Denote $U^{N}(t):=N(\widetilde{Y}(t)-Y(t))$, $t\in[0,T]$. Let Assumption 2 hold. Then $U^{N}$ has the following representation
$\;U^{N}(t)=\int_{0}^{t}\nabla\bar{f}(Y(s))U^{N}(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))U^{N}(s)\mathrm{d}W(s)+\sum_{i=1}^{4}I_{i}^{N}(t)+R_{U^{N}}(t).$ (4.2)
Here,
$\;I_{1}^{N}(t)=T\int_{0}^{t}\big{[}\alpha^{\top}(Ae)\nabla f(\widetilde{Y}(\kappa_{N}(s)))f(\widetilde{Y}(\kappa_{N}(s)))+F_{3}(\widetilde{Y}(\kappa_{N}(s)))+3F_{4}(\widetilde{Y}(\kappa_{N}(s)))\big{]}\mathrm{d}s,$
$\;I_{2}^{N}(t)=T\int_{0}^{t}\big{[}F_{1}(\widetilde{Y}(\kappa_{N}(s)))+3F_{2}(\widetilde{Y}(\kappa_{N}(s)))-\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))g(\widetilde{Y}(\kappa_{N}(s)))\big{]}\mathrm{d}W(s),$
$\;I_{3}^{N}(t)=T\int_{0}^{t}\big{[}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\bar{f}(\widetilde{Y}(\kappa_{N}(s)))+\frac{1}{2}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))\big{]}$
$\;\big{(}\lfloor\frac{Ns}{T}\rfloor-\frac{Ns}{T}\big{)}\mathrm{d}s,$
$\;I_{4}^{N}(t)=T\int_{0}^{t}G_{1}(\widetilde{Y}(\kappa_{N}(s)))(\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor)\mathrm{d}W(s)$
$\;\qquad\qquad+N\int_{0}^{t}G_{2}(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}W(s),$
where $G_{1}(y):=\nabla\bar{f}(y)g(y)-\nabla g(y)\bar{f}(y)-\frac{1}{2}\mathcal{D}^{2}g(y)(g(y),g(y))$, $G_{2}(y)=6F_{2}(y)-(\nabla g(y))^{2}g(y)-\mathcal{D}^{2}g(y)(g(y),g(y))$, $y\in\mathbb{R}^{d}$ and $R_{U^{N}}(t)$ is the remainder with $\sup_{t\in[0,T]}\mathbf{E}|R_{U^{N}}(t)|^{2}\leq Kh$.
Proof.
Subtracting $Y(t)$ from $\widetilde{Y}(t)$, we have
$U^{N}(t)=$ $\;N\int_{0}^{t}\big{[}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))-\bar{f}(Y(s))\big{]}\mathrm{d}s+N\int_{0}^{t}[g(\widetilde{Y}(\kappa_{N}(s)))-g(Y(s))\big{]}\mathrm{d}W(s)$
$\;+N\int_{0}^{t}\int_{\kappa_{N}(s)}^{s}\nabla g(\widetilde{Y}(\kappa_{N}(r)))g(\widetilde{Y}(\kappa_{N}(r)))\mathrm{d}W(r)\mathrm{d}W(s)+T\int_{0}^{t}\big{[}F_{1}(\widetilde{Y}(\kappa_{N}(s)))$
$\;+3F_{2}(\widetilde{Y}(\kappa_{N}(s)))\big{]}\mathrm{d}W(s)+6N\int_{0}^{t}\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}F_{2}(\widetilde{Y}(\kappa_{N}(u)))\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}W(s)$
$\;+T\int_{0}^{t}\big{[}\alpha^{\top}(Ae)\nabla f(\widetilde{Y}(\kappa_{N}(s)))f(\widetilde{Y}(\kappa_{N}(s)))+F_{3}(\widetilde{Y}(\kappa_{N}(s)))+3F_{4}\widetilde{Y}(\kappa_{N}(s))\big{]}\mathrm{d}s.$ (4.3)
The Taylor formula yields
$\;N\int_{0}^{t}\big{[}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))-\bar{f}(Y(s))\big{]}\mathrm{d}s$
$=$ $\;N\int_{0}^{t}\big{[}\bar{f}(\widetilde{Y}(s))-\bar{f}(Y(s))\big{]}\mathrm{d}s-N\int_{0}^{t}\big{[}\bar{f}(\widetilde{Y}(s))-\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\big{]}\mathrm{d}s$
$=$ $\;N\int_{0}^{t}\nabla\bar{f}(Y(s))(\widetilde{Y}(s)-Y(s))\mathrm{d}s+A_{1}^{N}(t)-N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}s$
$\;-\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)),\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}s+A_{2}^{N}(t),$ (4.4)
where
$A_{1}^{N}(t)=$ $\;N\int_{0}^{t}\int_{0}^{1}(1-\lambda)\mathcal{D}^{2}\bar{f}(Y(s)+\lambda(\widetilde{Y}(s)-Y(s)))(\widetilde{Y}(s)-Y(s),\widetilde{Y}(s)-Y(s))\mathrm{d}\lambda\mathrm{d}s,$
$A_{2}^{N}(t)=$ $\;-\frac{N}{2}\int_{0}^{t}\int_{0}^{1}(1-\lambda)^{2}\mathcal{D}^{3}\bar{f}(\widetilde{Y}(\kappa_{N}(s))+\lambda(\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s))))\big{(}\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)),$
$\;\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)),\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s))\big{)}\mathrm{d}\lambda\mathrm{d}s.$
Based on a standard argument, we obtain
$\sup_{t\in[0,T]}\|\widetilde{Y}(t)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T),~\|\widetilde{Y}(t)-\widetilde{Y}(s)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)|t-s|^{1/2},~t,s\in[0,T],$ (4.5)
which combined with Lemma 4.1 and (3.3) gives $\|A_{1}^{N}(t)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h$ and $\|A_{2}^{N}(t)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h^{1/2}$. By (4.1), we can write
$\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s))=$ $\;\int_{\kappa_{N}(s)}^{s}g(\widetilde{Y}(\kappa_{N}(r)))\mathrm{d}W(r)+\int_{\kappa_{N}(s)}^{s}\bar{f}(\widetilde{Y}(\kappa_{N}(r)))\mathrm{d}r$
$\;+\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\nabla g(\widetilde{Y}(\kappa_{N}(u)))g(\widetilde{Y}(\kappa_{N}(u)))\mathrm{d}W(u)\mathrm{d}W(r)+R^{1}_{\widetilde{Y}}(s),$ (4.6)
with $\|R^{1}_{\widetilde{Y}}(s)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h^{3/2}$. Thus, by (3.14) one has
$\;\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)),\widetilde{Y}(s)-\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}s$
$\;=\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))(W(s)-W(\kappa_{N}(s)))^{2}\mathrm{d}s+A_{3}^{N}(t)$
$\;=\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))(s-\kappa_{N}(s))\mathrm{d}s+\tilde{R}_{1}^{N}(t)+A_{3}^{N}(t),$ (4.7)
with $\tilde{R}_{1}^{N}(t):=N\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}s$ and $\|A_{3}^{N}(t)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h^{1/2}$, $p\geq 1$. Plugging (4.6) into the third term of (4.4) and using (4.7), we get
$\;N\int_{0}^{t}\big{[}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))-\bar{f}(Y(s))\big{]}\mathrm{d}s$
$=$ $\;\int_{0}^{t}\nabla\bar{f}(Y(s))U^{N}(s)\mathrm{d}s-N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}g(\widetilde{Y}(\kappa_{N}(r)))\mathrm{d}W(r)\mathrm{d}s$
$\;-N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(s-\kappa_{N}(s))\mathrm{d}s$
$\;-\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))(s-\kappa_{N}(s))\mathrm{d}s+\tilde{R}_{3}^{N}(t),$ (4.8)
where $\tilde{R}_{3}^{N}(t)=-\tilde{R}_{1}^{N}(t)-\tilde{R}_{2}^{N}(t)+A_{1}^{N}(t)+A_{2}^{N}(t)-A_{3}^{N}(t)-N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))R^{1}_{\widetilde{Y}}(s)\mathrm{d}s$ and $\tilde{R}_{2}^{N}(t):=N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\nabla g(\widetilde{Y}(\kappa_{N}(s)))g(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}s$.
We write $\tilde{R}_{2}^{N}(t)$ as $\tilde{R}_{2}^{N}(t)=N\sum\limits_{i=0}^{\lfloor\frac{t}{h}\rfloor}M_{i}(t)$ with
$M_{i}(t)=\int_{t_{i}}^{t_{i+1}\wedge t}\nabla\bar{f}(\widetilde{Y}(t_{i}))\nabla g(\widetilde{Y}(t_{i}))g(\widetilde{Y}(t_{i}))\int_{t_{i}}^{s}\int_{t_{i}}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}s.$
By the independence of increments of $W$ and the properties of the condition expectation, for any $j>i$, $\mathbf{E}\langle M_{i}(t),M_{j}(t)\rangle=\mathbf{E}\langle M_{i}(t),\mathbf{E}(M_{j}(t)|\mathcal{F}_{t_{j}})\rangle=0$. Accordingly,
$\mathbf{E}|\tilde{R}_{2}^{N}(t)|^{2}=N^{2}\mathbf{E}\sum_{i=0}^{\lfloor\frac{t}{h}\rfloor}|M_{i}(t)|^{2}\leq KN^{3}h^{4}\leq Kh\to 0,~\text{as}~N\to\infty.$ (4.9)
Similarly, one has $\mathbf{E}|\tilde{R}_{1}^{N}(t)|^{2}\leq Kh$. Based on previous estimates, we obtain $\mathbf{E}|\tilde{R}_{3}^{N}(t)|^{2}\leq Kh$. By the stochastic Fubini theorem,
$\;N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}g(\widetilde{Y}(\kappa_{N}(r)))\mathrm{d}W(r)\mathrm{d}s$
$=$ $\;N\int_{0}^{t}\int_{r}^{(\kappa_{N}(r)+T/N)\wedge t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))g(\widetilde{Y}(\kappa_{N}(s)))\mathrm{d}s\mathrm{d}W(r)$
$=$ $\;N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(r)))g(\widetilde{Y}(\kappa_{N}(r)))(\kappa_{N}(r)+\frac{T}{N}-r)\mathrm{d}W(r)+\tilde{R}_{4}^{N}(t),$
where $\tilde{R}_{4}^{N}(t):=N\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(r)))g(\widetilde{Y}(\kappa_{N}(r)))\big{(}(\kappa_{N}(r)+\frac{T}{N})\wedge t-(\kappa_{N}(r)+\frac{T}{N})\big{)}\mathrm{d}W(r)$. Clearly, $\tilde{R}_{4}^{N}(t):=N\int_{\kappa_{N}(t)}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(t)))g(\widetilde{Y}(\kappa_{N}(t)))\big{(}t-(\kappa_{N}(t)+\frac{T}{N})\big{)}\mathrm{d}W(r)$ due to the definition of $\kappa_{N}$. Thus, the Itô isometry formula yields $\mathbf{E}|\tilde{R}_{4}^{N}(t)|^{2}\leq Kh$. Thus, (4.8) becomes
$\;N\int_{0}^{t}\big{[}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))-\bar{f}(Y(s))\big{]}\mathrm{d}s$
$=$ $\;\int_{0}^{t}\nabla\bar{f}(Y(s))U^{N}(s)\mathrm{d}s-T\int_{0}^{t}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))g(\widetilde{Y}(\kappa_{N}(s)))\big{(}\lfloor\frac{Ns}{T}\rfloor-\frac{Ns}{T}+1\big{)}\mathrm{d}W(s)$
$\;-T\int_{0}^{t}\big{[}\nabla\bar{f}(\widetilde{Y}(\kappa_{N}(s)))\bar{f}(\widetilde{Y}(\kappa_{N}(s)))+\frac{1}{2}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))\big{]}$
$\;\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}\mathrm{d}s+\tilde{R}_{3}^{N}(t)-\tilde{R}_{4}^{N}(t).$ (4.10)
Using the Taylor formula, $N\int_{0}^{t}\big{[}g(\widetilde{Y}(\kappa_{N}(s)))-g(Y(s))\big{]}\mathrm{d}W(s)$ has a similar expansion as (4.4), where $\bar{f}$ and $\mathrm{d}s$ are replaced by $g$ and $\mathrm{d}W(s)$, respectively. Further, using (4.6) and (3.14), we obtain
$\;N\int_{0}^{t}\big{[}g(\widetilde{Y}(\kappa_{N}(s)))-g(Y(s))\big{]}\mathrm{d}W(s)$
$=$ $\;\int_{0}^{t}\nabla g({Y}(s))U^{N}(s)\mathrm{d}W(s)-N\int_{0}^{t}\nabla g(\widetilde{Y}(\kappa_{N}(s)))g(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}\mathrm{d}W(r)\mathrm{d}W(s)$
$\;-N\int_{0}^{t}\Big{[}(\nabla g(\widetilde{Y}(\kappa_{N}(s))))^{2}g(\widetilde{Y}(\kappa_{N}(s)))+\mathcal{D}^{2}g(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))\Big{]}$
$\;\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}W(s)-T\int_{0}^{t}\big{[}\nabla g(\widetilde{Y}(\kappa_{N}(s)))\bar{f}(\widetilde{Y}(\kappa_{N}(s)))$
$\;+\frac{1}{2}\mathcal{D}^{2}g(\widetilde{Y}(\kappa_{N}(s)))\big{(}g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s)))\big{)}\big{]}(\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor)\mathrm{d}W(s)+\tilde{R}_{5}^{N}(t)$ (4.11)
with $\mathbf{E}|\tilde{R}_{5}^{N}(t)|^{2}\leq Kh$.
Inserting (4.10) and (4.11) into (4.3) yields (4.2) with $R_{U^{N}}(t)=\tilde{R}_{3}^{N}(t)-\tilde{R}_{4}^{N}(t)+\tilde{R}_{5}^{N}(t)$. According to the previous estimates for $\tilde{R}_{i}(t)$, $i=3,4,5$, it holds that $\mathbf{E}|R_{U^{N}}(t)|^{2}\leq Kh$ for any $t\in[0,T]$. ∎
Lemma 4.3.
Let Assumption 2 hold. Then for any $t\in[0,T]$, $\lim\limits_{N\to\infty}\mathbf{E}|U^{N}(t)-\widetilde{U}^{N}(t)|^{2}=0$. Here, $\widetilde{U}^{N}$ is the strong solution of the following equation
$\widetilde{U}^{N}(t)=\int_{0}^{t}\nabla\bar{f}(Y(s))\widetilde{U}^{N}(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))\widetilde{U}^{N}(s)\mathrm{d}W(s)+\sum_{i=1}^{3}I_{i}(t)+I^{N}_{4}(t),~t\in[0,T],$ (4.12)
where
$I_{1}(t)$ $=T\int_{0}^{t}\big{[}\alpha^{\top}(Ae)\nabla f(Y(s))f(Y(s))+F_{3}(Y(s))+3F_{4}(Y(s))\big{]}\mathrm{d}s,$
$I_{2}(t)$ $=T\int_{0}^{t}\big{[}F_{1}(Y(s))+3F_{2}(Y(s))-\nabla\bar{f}(Y(s))g(Y(s))\big{]}\mathrm{d}W(s),$
$I_{3}(t)$ $=-\frac{T}{2}\int_{0}^{t}\big{[}\nabla\bar{f}(Y(s))\bar{f}(Y(s))+\frac{1}{2}\mathcal{D}^{2}\bar{f}(Y(s))(g(Y(s)),g(Y(s)))\big{]}\mathrm{d}s.$
Proof.
We first show that for any $t\in[0,T]$, $I_{i}^{N}(t)$ converges to $I_{i}(t)$ in $\|\cdot\|_{\mathbf{L}^{2}(\Omega)}$-norm for $i=1,2,3$. It follows from (3.4) and Lemma 3.8 that
$\sup_{s\in[0,T]}\|\widetilde{Y}(\kappa_{N}(s))-Y(s)\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h^{1/2},~\forall~p\geq 1.$ (4.13)
For any Lipschitz function $H$, using (4.13), the Hölder inequality and the Itô isometry formula, one has
$\lim_{N\to\infty}\mathbf{E}\Big{|}\int_{0}^{t}\big{(}H(\widetilde{Y}(\kappa_{N}(s)))-H(Y(s))\Big{)}\mathrm{d}s\Big{|}^{2}$ =0=0,= 0 , (4.14)
$\lim_{N\to\infty}\mathbf{E}\Big{|}\int_{0}^{t}\big{(}H(\widetilde{Y}(\kappa_{N}(s)))-H(Y(s))\Big{)}\mathrm{d}W(s)\Big{|}^{2}$ =0=0.= 0 .
Thus, it holds that $\lim\limits_{N\to\infty}\mathbf{E}|I^{N}_{i}(t)-I_{i}(t)|^{2}=0$ for $i=1,2$ and $t\in[0,T]$. Moreover, applying (4.14) and [6, Proposition 4.2], we obtain $\lim\limits_{N\to\infty}\mathbf{E}|I^{N}_{3}(t)-I_{3}(t)|^{2}=0$ for any $t\in[0,T]$.
Note that $\nabla\bar{f}$ and $\nabla g$ are bounded functions under Assumption 2. It follows from (4.2), (4.12), the Itô isometry formula and the Hölder inequality that
$\mathbf{E}|U^{N}(t)-\widetilde{U}^{N}(t)|^{2}\leq K\mathbf{E}\int_{0}^{t}|U^{N}(s)-\widetilde{U}^{N}(s)|^{2}\mathrm{d}s+Ka^{N}(t),$
where $a^{N}(t):=\mathbf{E}\sum_{i=1}^{3}|I_{i}^{N}(t)-I_{i}(t)|^{2}+\mathbf{E}|R_{U^{N}}(t)|^{2}$. The Gronwall inequality yields
$\mathbf{E}|U^{N}(t)-\widetilde{U}^{N}(t)|^{2}\leq a^{N}(t)+K\int_{0}^{t}a^{N}(s)\mathrm{d}s.$
The convergence of $I_{i}^{N}(t)$, $i=1,2,3$ and $\sup_{t\in[0,T]}\mathbf{E}|R_{U^{N}}(t)|\leq Kh$ (see Lemma 4.2) lead to $\lim\limits_{N\to\infty}a^{N}(t)=0$ for any $t\in[0,T]$. By (4.5), we arrive at $\sup_{t\in[0,T]}|a^{N}(t)|\leq K(p,T)$. The proof is complete due to the dominated convergence theorem. ∎
Lemma 4.4.
Let Assumption 2 hold. Then $I_{4}^{N}\overset{satbly}{\Longrightarrow}I_{4}$ as $\mathbf{C}([0,T];\mathbb{R}^{d})$-valued random variables as $N\to\infty$. Here, $I_{4}$ is the stochastic process defined by
$I_{4}(t)=\frac{T}{2}\int_{0}^{t}G_{1}(Y(s))\mathrm{d}W(s)+\frac{T}{\sqrt{12}}\int_{0}^{t}G_{1}(Y(s))\mathrm{d}\widetilde{W}_{1}(s)+\frac{T}{\sqrt{6}}\int_{0}^{t}G_{2}(Y(s))\mathrm{d}\widetilde{W}_{2}(s),~t\in[0,T],$
where $(\widetilde{W}_{1},\widetilde{W}_{2})$ is a two-dimensional standard Brownian motion and independent of $W$.
Proof.
The proof is mainly based on [9, Theorem 4-1]. Hereafter, denote by $\langle X,Y\rangle_{t}$, $t\in[0,T]$ the cross variation process between the real-valued semi-martingales $\{X(t)\}_{t\in[0,T]}$ and $\{Y(t)\}_{t\in[0,T]}$. Let $I_{4}^{N,i}$ denote the $i$th component of $I_{4}^{N}$ for $i=1,\ldots,d$. Recall that
$I_{4}^{N}(t)=$ $\;T\int_{0}^{t}G_{1}(\widetilde{Y}(\kappa_{N}(s)))(\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor)\mathrm{d}W(s)$
$\;+N\int_{0}^{t}G_{2}(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}W(s).$
It is not hard to see $\langle I^{N,i}_{4},W\rangle_{t}=B_{1}^{N}(t)+B_{2}^{N}(t)$ with
$B_{1}^{N}(t)=$ $\;T\int_{0}^{t}G_{1}^{i}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}\mathrm{d}s,$
$B_{2}^{N}(t)=$ $\;N\int_{0}^{t}G_{2}^{i}(\widetilde{Y}(\kappa_{N}(s)))\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\mathrm{d}s.$
Applying (4.14) and [6, Proposition 4.2], we arrive at $\lim\limits_{N\to\infty}\mathbf{E}|B_{1}^{N}(t)-\frac{T}{2}\int_{0}^{t}G_{1}^{i}(Y(s))\mathrm{d}s|^{2}=0$. Further, similar to the proof of (4.9), one has $\lim\limits_{N\to\infty}\mathbf{E}|B_{2}^{N}(t)|^{2}=0$. Thus, for any $t\in[0,T]$ and $i=1,\ldots,d$,
$\langle I^{N,i}_{4},W\rangle_{t}\to\frac{T}{2}\int_{0}^{t}G_{1}^{i}(Y(s))\mathrm{d}s~\text{in probability},~\forall~t\in[0,T],~i=1,\ldots,d.$ (4.15)
Next, we derive the limit of $\langle I_{4}^{N,i},I_{4}^{N,j}\rangle_{t}$. A direct computation leads to
$\;\langle I_{4}^{N,i},I_{4}^{N,j}\rangle_{t}=T^{2}\int_{0}^{t}G_{1}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{1}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}^{2}\mathrm{d}s$
$\;+N^{2}\int_{0}^{t}G_{2}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{2}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\big{)}^{2}\mathrm{d}s+B_{3}^{N}(t),$
where $B_{3}^{N}(t)=TN\int_{0}^{t}\big{[}G_{1}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{2}^{j}(\widetilde{Y}(\kappa_{N}(s)))+G_{1}^{j}(\widetilde{Y}(\kappa_{N}(s)))G_{2}^{i}(\widetilde{Y}(\kappa_{N}(s)))\big{]}\\ \int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}\mathrm{d}s$. It follows from the Itô formula and (3.14) that
$\;\Big{(}\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}\mathrm{d}W(u)\mathrm{d}W(r)\Big{)}^{2}=2\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}(W(u)-W(\kappa_{N}(s)))\mathrm{d}W(u)\big{(}W(r)-W(\kappa_{N}(s))\big{)}\mathrm{d}W(r)$
$\;+2\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}(W(u)-W(\kappa_{N}(s)))\mathrm{d}W(u)\mathrm{d}r+\frac{1}{2}(s-\kappa_{N}(s))^{2}.$
Combining the above relations, we have
$\langle I_{4}^{N,i},I_{4}^{N,j}\rangle_{t}=$ $\;T^{2}\int_{0}^{t}G_{1}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{1}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}^{2}\mathrm{d}s$
$\;+\frac{T^{2}}{2}\int_{0}^{t}G_{2}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{2}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}^{2}\mathrm{d}s+B_{3}^{N}(t)+B_{4}^{N}(t),$
where
$\;B_{4}^{N}(t)=2N^{2}\int_{0}^{t}G_{2}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{2}^{j}(\widetilde{Y}(\kappa_{N}(s)))\Big{\{}\int_{\kappa_{N}(s)}^{s}\Big{[}\int_{\kappa_{N}(s)}^{r}(W(u)-W(\kappa_{N}(s)))\mathrm{d}W(u)$
$\;\big{(}W(r)-W(\kappa_{N}(s))\big{)}\Big{]}\mathrm{d}W(r)+\int_{\kappa_{N}(s)}^{s}\int_{\kappa_{N}(s)}^{r}(W(u)-W(\kappa_{N}(s)))\mathrm{d}W(u)\mathrm{d}r\Big{\}}\mathrm{d}s.$
Similar to the proof of (4.9), one can show that $\mathbf{E}|B_{3}^{N}(t)|^{2}+\mathbf{E}|B_{4}^{N}(t)|^{2}\leq Kh\to 0$, as $N\to\infty$. Further, using [6, Proposition 4.2] yields $\int_{0}^{t}G_{l}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{l}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}^{2}\mathrm{d}s$ converges to $\frac{1}{3}\int_{0}^{t}G_{l}^{i}(Y(s))G_{l}^{j}(Y(s))\mathrm{d}s$ in $\mathbf{L}^{2}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$ for $l=1,2$. In this way, we obtain
$\langle I_{4}^{N,i},I_{4}^{N,j}\rangle_{t}\to\frac{T^{2}}{3}\int_{0}^{t}G_{1}^{i}(Y(s))G_{1}^{j}(Y(s))\mathrm{d}s+\frac{T^{2}}{6}\int_{0}^{t}G_{2}^{i}(Y(s))G_{2}^{j}(Y(s))\mathrm{d}s$ (4.16)
in probability for any $t\in[0,T]$ and $i,j=1,\ldots,d$.
Combining (4.15)-(4.16) and using [9, Theorem 4-1], we obtain $I_{4}^{N}\overset{stably}{\Longrightarrow}\Psi$ as $\mathbf{C}([0,T];\mathbb{R}^{d})$-valued random variables as $N\to\infty$, and
$\;\langle\Psi^{i},W\rangle_{t}=\frac{T}{2}\int_{0}^{t}G_{1}^{i}(Y(s))\mathrm{d}s,$ (4.17)
$\;\langle\Psi^{i},\Psi^{j}\rangle_{t}=\frac{T^{2}}{3}\int_{0}^{t}G_{1}^{i}(Y(s))G_{1}^{j}(Y(s))\mathrm{d}s+\frac{T^{2}}{6}\int_{0}^{t}G_{2}^{i}(Y(s))G_{2}^{j}(Y(s))\mathrm{d}s.$ (4.18)
Further, it follows from [9, Proposition 1-4] that $\Psi^{i}$, $i=1,\ldots,d$ can be represented as
$\Psi^{i}(t)=\int_{0}^{t}u^{i}(s)\mathrm{d}W(s)+\sum_{k=1}^{d}\int_{0}^{t}v^{i,k}(s)\mathrm{d}\widetilde{W}_{k}(s),$
where $\widetilde{W}=\big{(}\widetilde{W}_{1},\ldots,\widetilde{W}_{d}\big{)}$ is a $d$-dimensional standard Brownian motion and independent of $W$. By (4.17), we have $u^{i}(s)=\frac{T}{2}G_{1}^{i}(Y(s))$, $i=1,\ldots,d$, which along with (4.18) gives
$\sum_{k=1}^{d}v^{i,k}(s)v^{j,k}(s)=\frac{T^{2}}{12}G_{1}^{i}(Y(s))G_{1}^{j}(Y(s))+\frac{T^{2}}{6}G_{2}^{i}(Y(s))G_{2}^{j}(Y(s)).$
We can take $v^{i,k}(s)=0$, $i=1,\ldots,d$, $k=3,\ldots,d$, $v^{i,1}(s)=\frac{T}{\sqrt{12}}G_{1}^{i}(Y(s))$ and $v^{i,2}(s)=\frac{T}{\sqrt{6}}G_{2}^{i}(Y(s))$, $i=1,\ldots,d$. Thus the proof is complete by letting $I_{4}=\Psi$. ∎
With previous preparation, now we can give the limit distribution of $U^{N}(t)$.
Theorem 4.5.
[…]
$\;U(t)=\int_{0}^{t}\nabla\bar{f}(Y(s))U(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))U(s)\mathrm{d}W(s)$
$\;+T\int_{0}^{t}\big{[}\alpha^{\top}(Ae)\nabla f(Y(s))f(Y(s))+F_{3}(Y(s))+3F_{4}(Y(s))\big{]}\mathrm{d}s$
$\;-\frac{T}{2}\int_{0}^{t}\big{[}\nabla\bar{f}(Y(s))\bar{f}(Y(s))+\frac{1}{2}\mathcal{D}^{2}\bar{f}(Y(s))(g(Y(s)),g(Y(s)))\big{]}\mathrm{d}s$
$\;+T\int_{0}^{t}\big{[}F_{1}(Y(s))+3F_{2}(Y(s))-\nabla\bar{f}(Y(s))g(Y(s))\big{]}\mathrm{d}W(s)$
$\;+T\int_{0}^{t}\big{[}\nabla\bar{f}(Y(s))g(Y(s))-\nabla g(Y(s))\bar{f}(Y(s))-\frac{1}{2}\mathcal{D}^{2}g(Y(s))(g(Y(s)),g(Y(s)))\big{]}\mathrm{d}\Big{(}\frac{W(s)}{2}+\frac{\widetilde{W}_{1}(s)}{\sqrt{12}}\Big{)}$
$\;+\frac{T}{\sqrt{6}}\int_{0}^{t}\big{[}6F_{2}(Y(s))-(\nabla g(Y(s)))^{2}g(Y(s))-\mathcal{D}^{2}g(Y(s))(g(Y(s)),g(Y(s)))\big{]}\mathrm{d}\widetilde{W}_{2}(s).$ (4.19)
Here, $(\widetilde{W}_{1},\widetilde{W}_{2})$ is a two-dimensional standard Brownian motion and independent of $W$. In addition, $F_{i}$, $i=1,2,3,4$ are those defined in Lemma 3.4.
Proof.
By Proposition 2.2 and Lemma 4.4, $(I_{1},I_{2},I_{3},I_{4}^{N})\overset{stably}{\Longrightarrow}(I_{1},I_{2},I_{3},I_{4})$ in $\mathbf{C}([0,T];\mathbb{R}^{d})^{\otimes 4}$ as $N\to\infty$. Note that a continuous mapping maps a family of random variables converging stably in law into another family of random variables converging stably in law, which can be directly shown according to the definition of stable convergence. Thus, we obtain $\Phi^{N}:=I_{1}+I_{2}+I_{3}+I_{4}^{N}\overset{stably}{\Longrightarrow}\Phi:=I_{1}+I_{2}+I_{3}+I_{4}$ in $\mathbf{C}([0,T];\mathbb{R}^{d})$.
Note that the coefficients $\nabla\bar{f}$ and $\nabla g$ of (4.12) are bounded and $\mathcal{D}^{2}\bar{f},\mathcal{D}^{2}g\in\mathbf{F}$. In addition, a direct computation leads to
$\sup_{N\geq 1}\sup_{t\in[0,T]}\mathbf{E}|\Phi^{N}(t)|^{4}+\sup_{t\in[0,T]}\mathbf{E}|\Phi(t)|^{4}<\infty,$
$\sup_{N\geq 1}\mathbf{E}|\Phi^{N}(t)-\Phi^{N}(s)|^{2}+\mathbf{E}|\Phi(t)-\Phi(s)|^{2}\leq K|t-s|.$
[…]
$V(t)=\int_{0}^{t}\nabla\bar{f}(Y(s))V(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))V(s)\mathrm{d}W(s)+\sum_{i=1}^{4}I_{i}(t),~t\in[0,T].$
[…]
4.2. Properties of limit equation
In this part, we study the properties of $U(t)$ and reveal the relationship between $\mathbf{E}|U(T)|^{2}$ and the weak convergence of the SRK method (3.9). A direct computation yields
$(\nabla\bar{f})\bar{f}$ $=(\nabla f)f+\frac{1}{2}(\nabla f\nabla g)g+\frac{1}{2}\mathcal{D}^{2}g(f,g)+\frac{1}{4}\mathcal{D}^{2}g(g,(\nabla g)g)+\frac{1}{2}(\nabla g)^{2}f+\frac{1}{4}(\nabla g)^{3}g,$
$\mathcal{D}^{2}\bar{f}(g,g)$ $=\mathcal{D}^{2}f(g,g)+\frac{1}{2}(\nabla g)\mathcal{D}^{2}g(g,g)+\mathcal{D}^{2}g(g,(\nabla g)g)+\frac{1}{2}\mathcal{D}^{3}g(g,g,g),$
$(\nabla\bar{f})g$ $=(\nabla f)g+\frac{1}{2}\mathcal{D}^{2}g(g,g)+\frac{1}{2}(\nabla g)^{2}g.$
Plugging the above three relations into (4.19) and using the definitions of $F_{i}$, $i=1,2,3,4$ (see Lemma 3.4), we have
$U(t)=$ $\;\int_{0}^{t}\nabla\bar{f}(Y(s))U(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))U(s)\mathrm{d}W(s)+T\int_{0}^{t}\mathbf{H}_{1}(Y(s))\mathrm{d}s$
$\;+T\int_{0}^{t}\mathbf{H}_{2}(Y(s))\mathrm{d}W(s)+\frac{T}{\sqrt{12}}\int_{0}^{t}\big{[}\nabla f(Y(s))g(Y(s))-\nabla g(Y(s))f(Y(s))\big{]}\mathrm{d}\widetilde{W}_{1}(s)$
$\;+\frac{T}{\sqrt{6}}\int_{0}^{t}\mathbf{H}_{3}(Y(s))\mathrm{d}\widetilde{W}_{2}(s),$ (4.20)
where
$\;\mathbf{H}_{1}(y)=\big{(}\alpha^{\top}(Ae)-\frac{1}{2}\big{)}\nabla f(y)f(y)+\big{(}\alpha^{\top}(B(Be))-\frac{1}{4}\big{)}\nabla f(y)\nabla g(y)g(y)$
$\;+\frac{1}{2}\big{(}\alpha^{\top}(Be)^{2}-\frac{1}{2}\big{)}\mathcal{D}^{2}f(y)(g(y),g(y))+\beta^{\top}(A(Be))\nabla g(y)\nabla f(y)g(y)$
$\;+\frac{3}{2}\big{(}\beta^{\top}(B(Be)^{2})-\frac{1}{12}\big{)}\nabla g(y)\mathcal{D}^{2}g(y)(g(y),g(y))+\big{(}\beta^{\top}(B(Ae))-\frac{1}{4}\big{)}(\nabla g(y))^{2}f(y)$
$\;+3\big{(}\beta^{\top}(B(B(Be)))-\frac{1}{24}\big{)}(\nabla g(y))^{3}g(y)+\big{(}\beta^{\top}((Ae)*(Be))-\frac{1}{4}\big{)}\mathcal{D}^{2}g(y)(f(y),g(y))$
$\;+3\big{(}\beta^{\top}((Be)*(B(Be)))-\frac{1}{8}\big{)}\mathcal{D}^{2}g(y)(g(y),\nabla g(y)g(y))+\frac{1}{2}\big{(}\beta^{\top}(Be)^{3}-\frac{1}{4}\big{)}\mathcal{D}^{3}g(y)(g(y),g(y),g(y)),$
$\;\mathbf{H}_{2}(y)=\big{(}\alpha^{\top}(Be)-\frac{1}{2}\big{)}\nabla f(y)g(y)+\big{(}\beta^{\top}(Ae)-\frac{1}{2}\big{)}\nabla g(y)f(y)$
$\;+3\big{(}\beta^{\top}(B(Be))-\frac{1}{6}\big{)}(\nabla g(y))^{2}g(y)+\frac{3}{2}\big{(}\beta^{\top}(Be)^{2}-\frac{1}{3}\big{)}\mathcal{D}^{2}g(y)(g(y),g(y)),$
$\;\mathbf{H}_{3}(y)=6\big{(}\beta^{\top}(B(Be))-\frac{1}{6}\big{)}(\nabla g(y))^{2}g(y)+3\big{(}\beta^{\top}(Be)^{2}-\frac{1}{3}\big{)}\mathcal{D}^{2}g(y)(g(y),g(y)).$
Define
$\eta_{1}=$ $\;(\alpha^{\top}(Ae)-\frac{1}{2})^{2}+(\alpha^{\top}(B(Be))-\frac{1}{4})^{2}+(\alpha^{\top}(Be)^{2}-\frac{1}{2})^{2}+(\beta^{\top}(A(Be)))^{2}+(\beta^{\top}(B(Be)^{2})-\frac{1}{12})^{2}$
$\;+(\beta^{\top}(B(Ae))-\frac{1}{4})^{2}+(\beta^{\top}(B(B(Be)))-\frac{1}{24})^{2}+(\beta^{\top}((Ae)*(Be))-\frac{1}{4})^{2}$
$\;+(\beta^{\top}((Be)*(B(Be)))-\frac{1}{8})^{2}+(\beta^{\top}(Be)^{3}-\frac{1}{4})^{2}+(\alpha^{\top}(Be)-\frac{1}{2})^{2}+(\beta^{\top}(Ae)-\frac{1}{2})^{2}$
$\;+(\beta^{\top}(B(Be))-\frac{1}{6})^{2}+(\beta^{\top}(Be)^{2}-\frac{1}{3})^{2}.$ (4.21)
Note that $\int_{0}^{t}\nabla\bar{f}(Y(s))U(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))U(s)\mathrm{d}W(s)=\int_{0}^{t}\nabla f(Y(s))U(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))U(s)\circ\mathrm{d}W(s)$. Next, we have the following upper estimate for $\mathbf{E}|U(T)|^{2}$.
Theorem 4.6.
Let Assumption 2 hold and $T\geq 1$. Then there exists $L_{1}>0$ independent of $T$ such that
$\mathbf{E}|U(T)|^{2}\leq e^{L_{1}T}(1+\eta_{1})T^{3}.$
Proof.
In this proof, we denote by $L$ a generic constant independent of $T$, which may vary for each appearance. Note that under Assumption 2, every function from $\mathcal{H}$ grows at most linearly, and $\nabla\bar{f}$ and $\nabla g$ are bounded. The Itô formula yields
$\;\mathbf{E}|U(t)|^{2}=2\mathbf{E}\int_{0}^{t}\langle U(s),\nabla\bar{f}(Y(s))U(s)\rangle\mathrm{d}s+2T\mathbf{E}\int_{0}^{t}\langle U(s),\mathbf{H}_{1}(Y(s))\mathrm{d}s$
$\;+\mathbf{E}\int_{0}^{t}\big{|}\nabla g(Y(s))U(s)+T\mathbf{H}_{2}(Y(s))|^{2}\mathrm{d}s+\frac{T^{2}}{12}\mathbf{E}\int_{0}^{t}|\nabla f(Y(s))g(Y(s))-\nabla g(Y(s))f(Y(s))|^{2}\mathrm{d}s$
$\;+\frac{T^{2}}{6}\mathbf{E}\int_{0}^{t}|\mathbf{H}_{3}(Y(s))|^{2}\mathrm{d}s$
$\leq L\int_{0}^{t}\mathbf{E}|U(s)|^{2}\mathrm{d}s+LT^{2}\sum_{i=1}^{3}\int_{0}^{t}\mathbf{E}|\mathbf{H}_{i}(Y(s))|^{2}\mathrm{d}s+LT^{2}\int_{0}^{t}(1+\mathbf{E}|Y(s)|^{2})\mathrm{d}s.$
By the definitions of $\eta_{1}$ (see (4.21)) and $\mathbf{H}_{i}$, $i=1,2,3$, we have
$\mathbf{E}|U(t)|^{2}\leq$ $\;L\int_{0}^{t}\mathbf{E}|U(s)|^{2}\mathrm{d}s+LT^{2}\int_{0}^{t}(1+\eta_{1})(1+\mathbf{E}|Y(s)|^{2})\mathrm{d}s.$ (4.22)
Applying (3.3), one has $\sup_{s\in[0,T]}\mathbf{E}|Y(s)|^{2}\leq Le^{LT}$. This, combined with (4.22) and the Gronwall inequality, yields the desired result. ∎
Next, we show that the SRK method (3.9) is of weak order 222 provided $\eta_{1}=0$.
Lemma 4.7.
Let Assumptions 2 hold. Assume that $f\in\mathbf{C}^{6}(\mathbb{R}^{d})$ and $g\in\mathbf{C}^{7}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}f,\,\mathcal{D}^{7}g\in\mathbf{F}$. If $\eta_{1}=0$, then the SRK method (3.9) is of weak order 222, i.e., for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$,
$\sup_{n=0,1,\ldots,N}\big{|}\mathbf{E}\varphi(Y^{SRK}_{n})-\mathbf{E}\varphi(Y(t_{n}))\big{|}\leq Kh^{2}.$
Proof.
In this proof, $M=\mathcal{O}(h^{k})$, $k>0$, means that for any $p\geq 1$, $\|M\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+y^{\iota})h^{k}$, $\iota\geq 1$, i.e., $M$ is of strong order $k$ in terms of $h$. Consider the following $s_{0}$-stage SRK for (3.1).
$\begin{cases}Z_{n,i}=P_{n}+h\sum\limits_{j=1}^{s_{0}}a_{ij}f(Z_{n,j})+\Delta W_{n}\sum\limits_{j=1}^{s_{0}}b_{ij}g(Z_{n,j}),~i=1,\ldots,s_{0},\\ P_{n+1}=P_{n}+h\sum\limits_{i=1}^{s_{0}}\alpha_{i}f(Z_{n,i})+\Delta W_{n}\sum\limits_{i=1}^{s_{0}}\beta_{i}g(Z_{n,i}),~n=0,\ldots,N-1,\end{cases}$
with $P_{0}=Y_{0}$. In [20], for the scalar noise case, it has been shown that $\{P_{n}\}_{n=0}^{N}$ converges to $Y$ if conditions [20, (1)-(17)] holds. Note that [20, (1)-(17)] is equivalent to $\eta_{1}=0$, $\alpha^{\top}e=\beta^{\top}e=1$, and $\beta^{\top}(Be)=\frac{1}{2}$. Thus, under the assumptions of the theorem,
$\sup_{n=0,1,\ldots,N}\big{|}\mathbf{E}\varphi(P_{n})-\mathbf{E}\varphi(Y(t_{n}))\big{|}\leq Kh^{2}.$ (4.23)
By revisiting the proof of (3.12), the one-step approximation generating $\{P_{n}\}_{n=0}^{N}$ is
$P_{t,y}(t+h)=$ $\;y+\Delta Wg(y)+hf(y)+\frac{1}{2}\Delta W^{2}\nabla g(y)g(y)+\Delta WhF_{1}(y)+\Delta W^{3}F_{2}(y)$
$\;+h^{2}\alpha^{\top}(Ae)\nabla f(y)f(y)+\Delta W^{2}hF_{3}(y)+\Delta W^{4}F_{4}(y)+\mathcal{O}(h^{2.5}).$ (4.24)
Denote $\bar{\Delta}=P_{t,y}(t+h)-y$. Then for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$, the Taylor formula gives
$\varphi(P_{t,y}(t+h))=$ $\;\varphi(y)+\sum_{i=1}^{d}\frac{\partial\varphi(y)}{\partial y_{i}}\bar{\Delta}^{i}+\frac{1}{2}\sum_{i_{1},i_{2}=1}^{d}\frac{\partial^{2}\varphi(y)}{\partial y_{i_{1}}\partial y_{i_{2}}}\bar{\Delta}^{i_{1}}\bar{\Delta}^{i_{2}}$
$\;+\cdots+\frac{1}{5!}\sum_{i_{1},\ldots,i_{5}=1}^{d}\frac{\partial^{5}\varphi(y)}{\partial y_{i_{1}}\partial y_{i_{2}}\cdots\partial y_{i_{5}}}\bar{\Delta}^{i_{1}}\bar{\Delta}^{i_{2}}\cdots\bar{\Delta}^{i_{5}}+\mathcal{O}(h^{3}),$ (4.25)
where $\bar{\Delta}^{i}$ is the $i$th component of $\bar{\Delta}$. Denote $\bar{\bar{\Delta}}=Y^{SRK}_{t,y}(t+h)-y$. Then
$\bar{\bar{\Delta}}=$ $\;\Delta\widehat{W}g(y)+hf(y)+\frac{1}{2}\Delta\widehat{W}^{2}\nabla g(y)g(y)+\Delta\widehat{W}hF_{1}(y)+\Delta\widehat{W}^{3}F_{2}(y)$
$\;+h^{2}\alpha^{\top}(Ae)\nabla f(y)f(y)+\Delta\widehat{W}^{2}hF_{3}(y)+\Delta\widehat{W}^{4}F_{4}(y)+\mathcal{O}(h^{2.5})$
due to (3.12). Then $\varphi(Y^{SRK}_{t,y}(t+h))$ also has a similar expansion as (4.25):
$\varphi(Y^{SRK}_{t,y}(t+h))=$ $\;\varphi(y)+\sum_{i=1}^{d}\frac{\partial\varphi(y)}{\partial y_{i}}\bar{\bar{\Delta}}^{i}+\frac{1}{2}\sum_{i_{1},i_{2}=1}^{d}\frac{\partial^{2}\varphi(y)}{\partial y_{i_{1}}\partial y_{i_{2}}}\bar{\bar{\Delta}}^{i_{1}}\bar{\bar{\Delta}}^{i_{2}}$
$\;+\cdots+\frac{1}{5!}\sum_{i_{1},\ldots,i_{5}=1}^{d}\frac{\partial^{5}\varphi(y)}{\partial y_{i_{1}}\partial y_{i_{2}}\cdots\partial y_{i_{5}}}\bar{\bar{\Delta}}^{i_{1}}\bar{\bar{\Delta}}^{i_{2}}\cdots\bar{\bar{\Delta}}^{i_{5}}+\mathcal{O}(h^{3}).$ (4.26)
In order to show that $\{Y^{SRK}_{n}\}_{n=0}^{N}$ is of weak order 222, it suffices to prove that $|\mathbf{E}\varphi(Y^{SRK}_{t,y}(t+h))-\mathbf{E}\varphi(P_{t,y}(t+h))|\leq K(1+|y|^{\iota_{1}})h^{3}$ for some $\iota_{1}\geq 1$ in view of (4.23). Thus the proof is complete once we show that there is some $K(y)\in\mathbf{F}$ such that for $k=1,\dots,5$,
$\mathbf{E}\big{(}\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}-\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}\big{)}\leq K(y)h^{3},~i_{j}\in\{1,\ldots,d\},$ (4.27)
due to (4.25) and (4.26).
Note that among all terms in the final expansions of $\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}$ and $\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}$, the expectations of those terms with strong order 0.50.50.5, 1.51.51.5 and 2.52.52.5 will vanish due to symmetry of distributions of $\Delta\widehat{W}$ and $\Delta W$. Thus, (4.27) holds if the expectations of those terms of strong 111 and 222 in the expansions of $\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}$ is 333th order approximation for corresponding terms in the expansions of $\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}$, which is further implied by
$|\mathbf{E}\big{(}\Delta\widehat{W}^{2}-\Delta W^{2})|+|\mathbf{E}\big{(}\Delta\widehat{W}^{4}-\Delta W^{4})|\leq Kh^{3}.$ (4.28)
Using (3.8) and $\kappa\geq 3$, we finally obtain (4.28) and thus finish the proof. ∎
Combining Theorems 4.5, Lemma 4.7, and (4.20), we obtain the following result.
Theorem 4.8.
Let the assumptions of Lemma 4.7 hold. If $\eta_{1}=0$, then the SRK method (3.9) is of weak order 222, and the limit distribution of $N(Y^{SRK}_{N}-Y(T))$ as $N\to\infty$ takes the following unified form
$U(T)=$ $\;\int_{0}^{T}\nabla\bar{f}(Y(s))U(s)\mathrm{d}s+\int_{0}^{T}\nabla g(Y(s))U(s)\mathrm{d}W(s)$
$\;+\frac{T}{\sqrt{12}}\int_{0}^{T}\big{[}\nabla f(Y(s))g(Y(s))-\nabla g(Y(s))f(Y(s))\big{]}\mathrm{d}\widetilde{W}_{1}(s).$
Remark 4.9.
[…]
5. Asymptotic error distribution for additive noise
In this section, we investigate the asymptotic error distribution of a class of SRK methods applied to SDEs with additive noise by following the procedure of Framework 1. The study route and proof ideas in establishing the asymptotic error distribution are similar to the case of multiplicative noise. Thus, we focus on stating our results and omit most unnecessary proof details.
$\begin{cases}\mathrm{d}X(t)=f(X(t))\mathrm{d}t+\sigma\mathrm{d}\mathbf{W}(t),~t\in(0,T],\\ X(0)=X_{0}\in\mathbb{R}^{d},\end{cases}$ (5.1)
where $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is Lipschitz continuous, $X_{0}$ is non-random, and $\sigma=(\sigma_{1},\ldots,\sigma_{m})\in\mathbb{R}^{d\times m}$ is a constant matrix.
Consider the following SRK method:
$\begin{cases}Z_{n,i}=X^{SRK}_{n}+h\sum\limits_{j=1}^{s_{0}}\bar{a}_{ij}f(Z_{n,j})+\bar{b}_{i}\sigma\Delta\mathbf{W}_{n},~i=1,\ldots,s_{0},\\ X^{SRK}_{n+1}=X_{n}^{SRK}+h\sum\limits_{i=1}^{s_{0}}\bar{\alpha}_{i}f(Z_{n,i})+\sigma\Delta\mathbf{W}_{n},~n=0,\ldots,N-1,\end{cases}$ (5.2)
where $\bar{a}_{ij}$, $\bar{b}_{i}$ and $\bar{\alpha}_{i}$ are given constants, and $\Delta\mathbf{W}_{n}=\mathbf{W}(t_{n+1})-\mathbf{W}(t_{n})$. The coefficients of (5.2) can be represented by the Butcher tableau of the form   $\bar{A}$ $\bar{b}$ $\bar{\alpha}^{\top}$ , where $\bar{A}=(\bar{a}_{ij})$, $\bar{b}=(\bar{b}_{1},\ldots,\bar{b}_{s_{0}})^{\top}$ and $\bar{\alpha}=(\bar{\alpha}_{1},\ldots,\bar{\alpha}_{s_{0}})^{\top}$. The one-step approximation for (5.2) is
$X^{SRK}_{t,x}(t+h)=x+h\sum_{i=1}^{s_{0}}\bar{\alpha}_{i}f(Z_{i})+\sigma\Delta\mathbf{W},$
where $\Delta\mathbf{W}=\mathbf{W}(t+h)-\mathbf{W}(t)$ and $Z=(Z_{1}^{\top},\ldots,Z_{s_{0}}^{\top})^{\top}$ is determined by the following equation
$Z=(e\otimes I_{d})x+h(\bar{A}\otimes I_{d})F(Z)+(\bar{b}\otimes I_{d})\sigma\Delta\mathbf{W}$ (5.3)
with $F(Z)=(f(Z_{1})^{\top},\ldots,f(Z_{s_{0}})^{\top})^{\top}$.
Assumption 3.
Assume that $f\in\mathbf{C}^{4}_{b}(\mathbb{R}^{d})$ and $\bar{\alpha}^{\top}e=1$.
Under the above assumption, we can establish the solvability and convergence of (5.2).
Lemma 5.1.
Let Assumption 3 hold. Then there is $h^{\prime}_{1}>0$ such that for any $h\in(0,h^{\prime}_{1}]$, $x\in\mathbb{R}^{d}$, $\omega\in\Omega$, (5.3) is uniquely solvable. Moreover, there is $C^{\prime}_{0}>0$ such that
$|Z_{i}-x|\leq C^{\prime}_{0}(1+|x|)h+C^{\prime}_{0}|\Delta\mathbf{W}|,~h\in(0,h_{1}).$
Lemma 5.2.
Let Assumptions 3 hold. Then we have the following.
  • (1)
    For any $p\geq 1$, $\sup\limits_{n=0,1,\ldots,N}\mathbf{E}|X^{SRK_{n}}|^{p}\leq K(p,T)$.
  • (2)
    The SRK method is of strong order 111, $i.e.,$ for any $p\geq 1$,
    $\sup_{n=0,1,\ldots,N}\|X^{SRK}_{n}-X(t_{n})\|_{\mathbf{L}^{p}(\Omega)}\leq K(p,T)h.$
  • (3)
    If additionally assume that $\bar{\alpha}^{\top}(\bar{A}e)=\bar{\alpha}^{\top}\bar{b}^{2}=\bar{\alpha}^{\top}\bar{b}=\frac{1}{2}$, then the SRK method (3.9) is of weak order 222, i.e., for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$,
    $\sup_{n=0,1,\ldots,N}\big{|}\mathbf{E}\varphi(X^{SRK}_{n})-\mathbf{E}\varphi(X(t_{n}))\big{|}\leq Kh^{2}.$
Proof.
The proof of (1) is similar to that of Lemma (3.2) and thus is omitted.
The proof of (2) is similar to that of Theorem 3.3, and we give the sketch of proof. Following the bootstrap argument in the proof of Theorem 3.3, we have
$X^{SRK}_{t,x}=x+\sigma\Delta\mathbf{W}+hf(x)+R_{X^{SRK}}$
with $|\mathbf{E}R_{X^{SRK}}|\leq K(1+|x|^{2})h^{2}$ and $\|R_{X^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|x|^{2})h^{1.5}$. In addition, by the Itô Taylor expansion, one has
$X_{t,x}(t+h)=x+\sigma\Delta\mathbf{W}+hf(x)+R_{X}$
with $|\mathbf{E}R_{X}|\leq K(1+|x|^{2})h^{2}$ and $\|R_{X}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|x|^{2})h^{1.5}$. Then the second conclusion follows by applying Theorem 2.1.
Following the argument for proving $|\mathbf{E}\varphi(Y^{SRK}_{t,y}(t+h))-\mathbf{E}\varphi(P_{t,y}(t+h))|\leq K(1+|y|^{\iota_{1}})h^{3}$ in Lemma 4.7 (or using [16, Page 100,Theorem 2.1]), the third conclusion holds provided that for $k=1,\ldots,5$,
$\mathbf{E}\big{(}\prod_{j=1}^{k}\Delta_{1}^{i_{j}}-\prod_{j=1}^{k}\Delta_{2}^{i_{j}}\big{)}\leq K(x)h^{3},~i_{j}\in\{1,\ldots,d\},$ (5.4)
with $\Delta_{1}=X_{t,x}(t+h)-x$, $\Delta_{2}=X^{SRK}_{t,x}(t+h)-x$, and $K(x)\in\mathbf{F}$. A direct computation yields
$\Delta_{1}=$ $\;\sum_{k=1}^{m}\sigma_{k}\Delta W_{k}+f(x)h+\sum_{k=1}^{m}\nabla f(x)\sigma_{k}\int_{t}^{t+h}\int_{t}^{s}\mathrm{d}W_{k}(r)\mathrm{d}s+\frac{1}{2}h^{2}\big{[}\nabla f(x)f(x)$
$\;+\frac{1}{2}\sum_{k=1}^{m}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{k})\big{]}+\sum_{k,l=1}^{m}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{l})\int_{t}^{t+h}\int_{t}^{s}\int_{t}^{r}\mathrm{d}W_{l}(u)\mathrm{d}W_{k}(r)\mathrm{d}s+\tilde{R}_{X},$
and
$\Delta_{2}=$ $\;\sum_{k=1}^{m}\sigma_{k}\Delta W_{k}+f(x)h+(\bar{\alpha}^{\top}\bar{b})h\sum_{k=1}^{m}\nabla f(x)\sigma_{k}\Delta W_{k}+(\bar{\alpha}^{\top}\bar{b}^{2})\frac{h}{2}\sum_{k,l=1}^{m}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{l})\Delta W_{k}\Delta W_{l}$
$\;+(\bar{\alpha}^{\top}(\bar{A}e))h^{2}\nabla f(x)f(x)+\tilde{R}_{X^{SRK}}$ (5.5)
with $|\mathbf{E}\tilde{R}_{X}|+|\mathbf{E}\tilde{R}_{X^{SRK}}|\leq K(1+|x|^{4})h^{3}$ and $\|\tilde{R}_{X}\|_{\mathbf{L}^{p}(\Omega)}+\|\tilde{R}_{X^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|x|^{3})h^{2.5}$. Using the assumptions on $\bar{A}$, $\bar{\alpha}$ and $\bar{b}$, one can finally prove (5.4). ∎
Next, we give the appurtenant numerical method for (5.2). According to (5.5) and (3.14), we can write $X^{SRK}_{t,x}(t+h)$ as
$X^{SRK}_{t,x}(t+h)=$ $\;x+\sigma\Delta\mathbf{W}+f(x)h+(\bar{\alpha}^{\top}\bar{b})h\nabla f(x)\sigma\Delta\mathbf{W}+(\bar{\alpha}^{\top}\bar{b}^{2})\frac{h^{2}}{2}\sum_{k=1}^{m}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{k})$
$\;+(\bar{\alpha}^{\top}(\bar{A}e))h^{2}\nabla f(x)f(x)+\bar{R}_{X^{SRK}},$
where $\bar{R}_{X^{SRK}}=\tilde{R}_{X^{SRK}}+(\bar{\alpha}^{\top}\bar{b}^{2})\frac{h}{2}\sum\limits_{k\neq l}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{l})\Delta W_{k}\Delta W_{l}+(\bar{\alpha}^{\top}\bar{b}^{2})h\sum_{k=1}^{m}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{k})\\ \cdot\int_{t}^{t+h}\int_{t}^{s}\mathrm{d}W_{k}(r)\mathrm{d}W_{k}(s)$. Thus, we immediately obtain $|\mathbf{E}\bar{R}_{X^{SRK}}|\leq K(1+|x|^{4})h^{3}$ and $\|\bar{R}_{X^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|x|^{3})h^{2}$ for $p\geq 1$. Then define the one-step approximation
$\widetilde{X}_{t,x}(t+h)=$ $\;x+\sigma\Delta\mathbf{W}+f(x)h+(\bar{\alpha}^{\top}\bar{b})h\nabla f(x)\sigma\Delta\mathbf{W}+(\bar{\alpha}^{\top}\bar{b}^{2})\frac{h^{2}}{2}\sum_{k=1}^{m}\mathcal{D}^{2}f(x)(\sigma_{k},\sigma_{k})$
$\;+(\bar{\alpha}^{\top}(\bar{A}e))h^{2}\nabla f(x)f(x).$
Further, we define $\widetilde{X}_{0}=X_{0}$ and $\widetilde{X}_{n+1}=\widetilde{X}_{t_{n},\widetilde{X}_{n}}(t_{n+1})$, $n=0,\ldots,N-1$. In what follows, we show that $\{\widetilde{X}_{n}\}_{n=0}^{N}$ is the appurtenant numerical method for (5.2).
Lemma 5.3.
Let Assumption 3 hold. Then $\{\widetilde{X}_{n}\}_{n=0}^{N}$ is of strong order 1.51.51.5 and 111 when approximating $\{X^{SRK}_{n}\}_{n=0}^{N}$ and $\{X(t)\}_{t\in[0,T]}$, respectively. More precisely, we have that for any $p\geq 1$,
$\sup_{n=0,1,\ldots,N}\|\widetilde{X}_{n}-X^{SRK}_{n}\|_{\mathbf{L}^{p}(\Omega)}\leq Kh^{1.5},$
$\sup_{n=0,1,\ldots,N}\|\widetilde{X}_{n}-X(t_{n})\|_{\mathbf{L}^{p}(\Omega)}\leq Kh.$
Proof.
According to the definition of $\widetilde{X}_{t,x}(t+h)$, it holds that
$\big{|}\mathbf{E}\big{(}\widetilde{X}_{t,x}(t+h)-X^{SRK}_{t,x}(t+h)\big{)}\big{|}\leq K(1+|x|^{4})h^{3},$
$\|\widetilde{X}_{t,x}(t+h)-X^{SRK}_{t,x}(t+h)\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|x|^{3})h^{2}.$
Further, it is not hard to verify that conditions of Theorem 2.1 are satisfied with $p_{1}=3$ and $p_{2}=2$. Thus, $\sup\limits_{n=0,1,\ldots,N}\|\widetilde{X}_{n}-X^{SRK}_{n}\|_{\mathbf{L}^{p}(\Omega)}\leq Kh^{1.5}$. This, combined with Lemma 5.2(2) and the Minkowski inequality, completes the proof. ∎
Next, we introduce the continuous version of $\{\widetilde{X}_{n}\}_{n=0}^{N}$.
$\;\widetilde{X}(t)=X_{0}+\sigma\mathbf{W}(t)+\int_{0}^{t}f(\widetilde{X}(\kappa_{N}(s)))\mathrm{d}s+(\bar{\alpha}^{\top}\bar{b})h\int_{0}^{t}\nabla f(\widetilde{X}(\kappa_{N}(s)))\sigma\mathrm{d}\mathbf{W}(s)$
$\;+h\int_{0}^{t}\Big{[}(\bar{\alpha}^{\top}(\bar{A}e)\nabla f(\widetilde{X}(\kappa_{N}(s)))f(\widetilde{X}(\kappa_{N}(s)))+\frac{1}{2}(\bar{\alpha}^{\top}\bar{b}^{2})\sum_{k=1}^{m}\mathcal{D}^{2}f(\widetilde{X}(\kappa_{N}(s)))(\sigma_{k},\sigma_{k})\Big{]}\mathrm{d}s.$
Define $V^{N}(t)=N(\widetilde{X}(t)-X(t))$, $t\in[0,T]$. As a counterpart of Lemma 4.2, we give the equation satisfied by $V^{N}$.
Lemma 5.4.
Let Assumptions 3 hold. Then $V^{N}$ has the following representation
$\;V^{N}(t)=\int_{0}^{t}\nabla f(X(s))V^{N}(s)\mathrm{d}s+\sum_{i=1}^{4}J_{i}^{N}(t)+R_{V^{N}}(t),$
where
$\;J_{1}^{N}(t)=-T\int_{0}^{t}\nabla f(\widetilde{X}(\kappa_{N}(s)))(\lfloor\frac{Ns}{T}\rfloor-\frac{Ns}{T}+1)\sigma\mathrm{d}\mathbf{W}(s),$
$\;J_{2}^{N}(t)=-T\int_{0}^{t}\big{[}\nabla f(\widetilde{X}(\kappa_{N}(s)))f(\widetilde{X}(\kappa_{N}(s)))+\frac{1}{2}\sum_{k=1}^{m}\mathcal{D}^{2}f(\widetilde{X}(\kappa_{N}(s)))(\sigma_{k},\sigma_{k})\big{]}(\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor)\mathrm{d}s,$
$\;J_{3}^{N}(t)=(\bar{\alpha}^{\top}\bar{b})T\int_{0}^{t}\nabla f(\widetilde{X}(\kappa_{N}(s)))\sigma\mathrm{d}\mathbf{W}(s),$
$\;J_{4}^{N}(t)=T\int_{0}^{t}\Big{[}(\bar{\alpha}^{\top}(\bar{A}e)\nabla f(\widetilde{X}(\kappa_{N}(s)))f(\widetilde{X}(\kappa_{N}(s)))+\frac{1}{2}(\bar{\alpha}^{\top}\bar{b}^{2}))\sum_{k=1}^{m}\mathcal{D}^{2}f(\widetilde{X}(\kappa_{N}(s)))(\sigma_{k},\sigma_{k})\Big{]}\mathrm{d}s$
and the remainder $R_{V^{N}}$ satisfies $\sup_{t\in[0,T]}\|R_{V^{N}}(t)\|_{\mathbf{L}^{2}(\Omega)}\leq Kh^{0.5}$.
The proof of Lemma 5.4 is similar to that of Lemma 4.2 and thus is omitted. Now we present the asymptotic error distribution of (5.2).
Theorem 5.5.
[…]
$V(t)=$ $\;\int_{0}^{t}\nabla f(X(s))V(s)\mathrm{d}s+\big{(}\bar{\alpha}^{\top}(\bar{A}e)-\frac{1}{2}\big{)}T\int_{0}^{t}\nabla f(X(s))f(X(s))\mathrm{d}s$
$\;+\frac{1}{2}\big{(}\bar{\alpha}^{\top}\bar{b}^{2}-\frac{1}{2}\big{)}T\sum_{k=1}^{m}\int_{0}^{t}\mathcal{D}^{2}f(X(s))(\sigma_{k},\sigma_{k})\mathrm{d}s+\big{(}\bar{\alpha}^{\top}\bar{b}-\frac{1}{2}\big{)}T\int_{0}^{t}\nabla f(X(s))\sigma\mathrm{d}\mathbf{W}(s)$
$\;-\frac{T}{\sqrt{12}}\int_{0}^{t}\nabla f(X(s))\sigma\mathrm{d}\widetilde{\mathbf{W}}(s).$
Here, $\widetilde{\mathbf{W}}$ is an $m$-dimensional standard Brownian motion independent of $\mathbf{W}$.
Proof.
It is not hard to verify that for any $t\in[0,T]$ and $p\geq 1$, $J_{3}^{N}(t)$ converges to $(\bar{\alpha}^{\top}\bar{b})T\int_{0}^{t}\nabla f(X(s))\sigma\mathrm{d}\mathbf{W}(s)$ in $\mathbf{L}^{p}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$, and $J_{4}^{N}(t)$ converges to
$T\int_{0}^{t}\Big{[}(\bar{\alpha}^{\top}(\bar{A}e))\nabla f(X(s))f(X(s))+\frac{1}{2}(\bar{\alpha}^{\top}\bar{b}^{2})\sum_{k=1}^{m}\mathcal{D}^{2}f(X(s))(\sigma_{k},\sigma_{k})\Big{]}\mathrm{d}s$
in $\mathbf{L}^{p}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$. Similar to the argument for $\|I_{3}^{N}(t)-I_{3}(t)\|_{\mathbf{L}^{p}(\Omega)}\to 0$ in the proof of Lemma 4.3, applying [6, Proposition 4.2] yields that $J_{2}^{N}(t)$ converges to
$-\frac{T}{2}\int_{0}^{t}\big{[}\nabla f(X(s))f(X(s))+\frac{1}{2}\sum_{k=1}^{m}\mathcal{D}^{2}f(X(s))(\sigma_{k},\sigma_{k})\big{]}\mathrm{d}s$
in $\mathbf{L}^{p}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$.
Further, following the ideas for proving Lemma 4.4, one can show that $J_{1}^{N}\overset{stably}{\Longrightarrow}J_{1}$ as $\mathbf{C}([0,T];\mathbb{R}^{d})$-valued random variables as $N\to\infty$, where
$J_{1}(t):=-\frac{T}{2}\int_{0}^{t}\nabla f(X(s))\sigma\mathrm{d}\mathbf{W}(s)-\frac{T}{\sqrt{12}}\int_{0}^{t}\nabla f(X(s))\sigma\mathrm{d}\widetilde{\mathbf{W}}(s),~t\in[0,T].$
Following the arguments for proving Lemma 4.3 and Theorem 4.5, we finally complete the proof. ∎
The following result gives the upper estimate of $\mathbf{E}|V(T)|^{2}$.
Theorem 5.6.
Let Assumption 3 hold and $T\geq 1$. Then there is $L_{2}>0$ independent of $T$ such that
$\mathbf{E}|V(T)|^{2}\leq e^{L_{2}T}(1+\eta_{2})T^{3},$
where $\eta_{2}=(\bar{\alpha}^{\top}(\bar{A}e)-\frac{1}{2})^{2}+(\bar{\alpha}^{\top}\bar{b}^{2}-\frac{1}{2})^{2}+(\bar{\alpha}^{\top}\bar{b}-\frac{1}{2})^{2}$.
The proof of Theorem 5.6 is similar to that of Theorem 4.6 and thus is omitted. Combining Theorem 5.5 and Lemma 5.2(3), we have the following facts.
Theorem 5.7.
Let Assumption 3 hold. If $\eta_{2}=0$, then the SRK method (5.2) is of weak order 222, and the limit distribution of $N(X^{SRK}_{N}-X(T))$ as $N\to\infty$ takes the following unified form
$V(T)=\int_{0}^{T}\nabla f(X(s))V(s)\mathrm{d}s-\frac{T}{\sqrt{12}}\int_{0}^{T}\nabla f(X(s))\sigma\mathrm{d}\widetilde{\mathbf{W}}(s).$
Remark 5.8.
As is mentioned in Remark 4.9, Theorems 5.5 and 5.6 indicate that $\eta_{2}$ is the key parameter reflecting the growth rate of the mean-square error of the SRK method (5.2). In addition, we obtain the following inferences: (1) For the SRK methods (5.2) of strong order 111, the smaller $\eta_{2}$ is, the smaller the mean-square error is after a long time. (2) Among the SRK methods (5.2) of strong order 111, those of weak order 222, i.e., $\eta_{2}=0$, have the smallest mean-square errors after a long time.
6. Numerical experiments
In this section, we perform numerical experiments to verify our theoretical analysis. For convenience, in this section, we consider SDEs with additive noise. We verify the convergence in distribution of $N(X^{SRK}_{N}-X(T))$ in Example 6.1, and verify the inferences of Remark 5.8 in Example 6.2 by presenting the evolution of the mean-square error of numerical methods w.r.t. time.
Example 6.1.
Consider the following SDE:
$\begin{cases}\mathrm{d}X(t)=(-10X(t)+\sin(X(t)))\mathrm{d}t+\sigma_{1}\mathrm{d}W_{1}(t),~t\in(0,T],\\ X(0)=x_{0}\in\mathbb{R},\end{cases}$ (6.1)
where $W_{1}$ is a one-dimensional standard Brownian motion and $\sigma_{1}\in\mathbb{R}$ is a given constant.
In this experiment, we consider the trapezoid method for (6.1), which reads
$X_{n+1}=X_{n}+\frac{1}{2}h(f(X_{n})+f(X_{n+1}))+\sigma_{1}(W_{1}(t_{n+1})-W_{1}(t_{n})),~n=0,\ldots,N-1$ (6.2)
[…]
$V(T)=\int_{0}^{T}f^{\prime}(X(s))V(s)\mathrm{d}s-\frac{T}{\sqrt{12}}\int_{0}^{T}f^{\prime}(X(s))\sigma_{1}\mathrm{d}\widetilde{W}_{1}(s),$
where $\widetilde{W}_{1}$ is a one-dimensional standard Brownian motion and independent of $W_{1}$.
[…]
$V_{n+1}=V_{n}+hf^{\prime}(X_{n})V_{n}-\frac{T}{\sqrt{12}}f^{\prime}(X_{n})\sigma_{1}(\widetilde{W}_{1}(t_{n+1})-\widetilde{W}_{1}(t_{n})),~n=0,\ldots,N-1$
with $V_{0}=0$. The expectations of $\varphi(N(X_{N}-X(T)))$ and $\varphi(V(T))$ are realized by using the Monte–Carlo method with 150000150000150000 sample paths. Table 2 displays the error $\big{|}\mathbf{E}\varphi(N(X_{N}-X(T)))-\mathbf{E}\varphi(V(T))\big{|}$ for different step-sizes and test functions $\varphi$. We observe that this error tends to 0 as $h$ decreases. This verifies the conclusion of Theorem 5.5.
Table 2. Error between $\mathbf{E}\varphi(N(X_{N}-X(T)))$ and $\mathbf{E}\varphi(V(T))$ for different step-sizes and $\varphi$ for (6.1).
$\varphi$ $h$ error $2^{-3}$ $2^{-4}$ $2^{-5}$ $2^{-6}$ $2^{-6}$ $2^{-7}$
$\sin(x)$ 5.4543E-2 2.5259E-2 1.2327E-2 6.8793226E-3 3.0687E-3 1.3684E-3
$\sin(x^{3})$ 3.7141E-3 1.7050E-3 8.3550E-4 4.2827E-4 2.0214E-4 1.0498E-4
Example 6.2.
Consider the following one-dimensional SDE:
$\begin{cases}\mathrm{d}X(t)=(X(t)+\ln(1+X(t)^{2}))\mathrm{d}t+\sigma_{1}\mathrm{d}W_{1}(t)+\sigma_{2}\mathrm{d}W_{2}(t),\\ X(0)=x_{0}\in\mathbb{R},\end{cases}$ (6.3)
where $W_{1}$ and $W_{2}$ are two independent one-dimensional standard Brownian motions, and $\sigma_{1},\sigma_{2}\in\mathbb{R}$ are given constants.
In this experiment, we consider four numerical methods for approximating (6.3): the trapezoid method, the $\theta$-method with $\theta=\frac{1}{2}$, $\frac{\sqrt{2}}{2}$, 111. The $\theta$-method is a 111-stage Runge–Kutta method taking of form (5.2), whose Butcher tableau is $\theta$ $\theta$ 111 . In addition, the $\frac{1}{2}$-method and 111-method are indeed the midpoint method and the implicit Euler method, respectively. It is not difficult to compute $\eta_{2}$ for these four methods.
Table 3. Values of $\eta_{2}$ for four methods
trapezoid method midpoint method $\frac{\sqrt{2}}{2}$-method implicit Euler method
$\eta_{2}$ 0 $\frac{1}{16}=0.0625$ $\frac{3-2\sqrt{2}}{2}\approx 0.08579$ $\frac{3}{4}$
First, we test the mean-square convergence orders for the above four methods. We set $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and $T=1$. These methods are realized by applying the fixed point iteration with a tolerance error $10^{-10}$. The exact solution $X(T)$ are approximated by the trapezoid method with the small step-size $h=2^{-16}$. The expectation is approximately obtained based on the Monte–Carlo method with 500050005000 sample paths. It is observed in Figure 1 and Table 4 that all these four methods have first order mean-square convergence. The mean-square error between the numerical solution of the trapezoid method and $X(T)$ is smaller than that between the numerical solution of the midpoint method and $X(T)$, although they are quite close.
Figure 1. Mean square errors for the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method applied to (6.3) in the log-log scale for five different step-sizes $h=2^{-4},2^{-5},2^{-6},2^{-7},2^{-8}$.
Table 4. Mean square errors for the trapezoid method and midpoint method for (6.3) for different step-sizes.
$h=2^{-4}$ $h=2^{-5}$ $h=2^{-6}$ $h=2^{-7}$ $h=2^{-8}$
trapezoid method 0.12435 0.062660 0.031007 0.015376 0.0078069
midpoint method 0.12542 0.063027 0.031137 0.015436 0.0078372
Figure 2. $|\mathbf{E}\exp(-X_{N})-\mathbf{E}\exp(-X(T))|$ for the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method applied to (6.3) in the log-log scale for five different step-sizes $h=2^{-4},2^{-5},2^{-6},2^{-7},2^{-8}$
Figure 3. Evolution of mean-square errors of the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method applied to (6.3) with $h=0.01$.
Second, we test the weak convergence orders of these methods. We take $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and $T=2^{-2}$. These methods are realized by applying the fixed point iteration with a tolerance error $10^{-10}$. The exact solution $X(T)$ are approximated by the trapezoid method with the small step-size $h=2^{-16}$. The expectation is approximately obtained based on the Monte–Carlo method with 200000200000200000 sample paths. Figure 2 indicates that the weak order of the trapezoid method is 222 and the weak order of the other three methods is 111.
Finally, we perform numerical experiments to present the evolution of mean-square errors $(\mathbf{E}|X_{n}-X(t_{n})|^{2})^{1/2}$ of these four methods w.r.t. the time $t$. The exact solution is approximated by the trapezoid method with the step-size $h=10^{-4}$. We set $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and use 500050005000 sample paths to approximate the expectation. As is shown in Figure 3, the trapezoid method has the smallest mean-square error after a long time. Although the mean-square error of the trapezoid method is very close to that of the midpoint method over a period of short time, the difference between of them becomes considerably large after a long time. From Table 3, we know that $\eta_{2}=0$ for the trapezoid method, which is not the case for other three methods. This verifies the second inference of Remark 5.8. We also observe that the mean-square error of the $\frac{\sqrt{2}}{2}$-method is much smaller than that of the implicit Euler method for large time $T$. In addition, Table 3 shows that $\eta_{2}$ of the $\frac{\sqrt{2}}{2}$-method is smaller than that of the implicit Euler method, which verifies the first inference of Remark 5.8
7. Conclusion
In this work, we propose a framework to establish the asymptotic error distribution of diffusion-implicit or fully implicit numerical methods applied to SDEs. Based on the framework, we gives the asymptotic error distribution of the SRK methods of strong order 111 and identify the key parameter reflecting the growth rate of the mean-square error of the SRK methods by studying the properties of the limit distribution. We infer that among the SRK methods of strong order 111, those of weak order 222 have the smallest mean-square errors after a long time. These results suggest that in numerical simulations for SDEs, methods with high weak order should be prioritized among those sharing the same strong convergence order.
We would like to point out that Framework 1 also applies to numerical methods with strong order no less than 1.51.51.5 and SDEs driven by multidimensional Brownian motions for multiplicative noise case. Of course, more extensive computation is indispensable to achieve these generalizations. For example, for the case of multidimensional multiplicative noise, the SRK method only including the (truncated) Brownian increments is strong order 0.50.50.5 attainable; see [2, Theorem 4.1]. In this case, the normalized error process should be $\sqrt{N}(Y^{SRK}_{N}-Y(T))$. We anticipate that deriving the asymptotic error distribution of such an SRK method with strong order 0.50.50.5 would not require substantial computational effort. For the case of multidimensional multiplicative noise, to make the SRK methods attain strong order 111, the SRK methods should include multiple stochastic integrals or the underlying SDEs have commutative noise; see [2]. We will be devoted to solving these problems in the future work.
Appendix
Proof of Lemma 3.1
Proof.
Denote $\Phi_{y}(Z)=(e\otimes I_{d})y+h(A\otimes I_{d})F(Z)+\Delta\widehat{W}(B\otimes I_{d})G(Z)$. Under Assumption 1, $F$ and $Z$ are Lipschitz continuous, which gives that there is $K_{1}>0$ such that
$|\Phi_{y}(Z^{1})-\Phi_{y}(Z^{2})|\leq K_{1}(h+|\Delta\widehat{W}|)|Z^{1}-Z^{2}|.$
Further, there is $h_{2}>0$ such that for any $h\in(0,h_{2}]$ and $\omega\in\Omega$, $K_{1}(h+|\Delta\widehat{W}|)<1$. By the contraction mapping principle, for any $h\in(0,h_{2}]$, $\omega\in\Omega$ and $y\in\mathbb{R}^{d}$, $\Phi_{y}(Z)=Z$ has a unique solution $Z^{*}$, and $Z^{*}(\omega)=\lim\limits_{n\to\infty}Z^{n}(\omega)$, where $Z^{n+1}(\omega)=\Phi_{y}(Z^{n}(\omega))$ and $Z^{0}(\omega)=(e\otimes I_{d})y$.
Further, there is $K_{2}>0$ such that
$\;|\Phi_{y}(Z)-(e\otimes I_{d})y|\leq K(1+|Z|)(h+|\Delta\widehat{W}|)$
$\leq$ $\;K_{2}|Z-(e\otimes I_{d})y|(h+|\Delta\widehat{W}|)+K_{2}(1+|y|)(h+|\Delta\widehat{W}|).$
Choose $h_{1}\leq h_{2}$ such that for any $h\in(0,h_{1}]$, $K_{2}(h+|\Delta\widehat{W}|)<\frac{1}{2}$. Thus, for any $h\in(0,h_{1}]$, if $|Z-(e\otimes I_{d})y|\leq 2K_{2}(1+|y|)(h+|\Delta\widehat{W}|)$, then
$|\Phi_{y}(Z)-(e\otimes I_{d})y|\leq 2K_{2}(1+|y|)(h+|\Delta\widehat{W}|),$
which along with the definition of $Z^{n}$ yields $|Z^{n}-(e\otimes I_{d})y|\leq 2K_{2}(1+|y|)(h+|\Delta\widehat{W}|)$. Passing to the limit produces $|Z^{*}-(e\otimes I_{d})y|\leq C_{0}(1+|y|)(h+|\Delta\widehat{W}|)$ with $C_{0}:=2K_{2}$. ∎
Proof of Lemma 3.2
Proof.
Note that $h+|\Delta\widehat{W}|\leq K$ for any $h\leq h_{1}\wedge 1$, which combined with Lemma 3.1 gives that $|Z_{i}|\leq K|y|+K(h+|\Delta\widehat{W}|)$. Thus, by (3.10) and the Lipschitz property of $f$ and $g$, $|Y^{SRK}_{t,x}(t+h)-y|\leq K(h+|\Delta\widehat{W}|)\sum_{i=1}^{s_{0}}(1+|Z_{i}|)\leq K(1+|y|)(h+|\Delta\widehat{W}|)=M(\Delta\widehat{W})(1+|y|)h^{1/2}$ with $M(\Delta\widehat{W}):=K(h^{1/2}+|\Delta\widehat{W}|h^{-1/2})$. It is easy to see that $\mathbf{E}|M(\Delta\widehat{W})|^{p}\leq K(p)$ for any $p\geq 1$. Further, from Lemma 3.1 we obtain
$\big{|}\mathbf{E}\big{[}\Delta\widehat{W}\sum_{i=1}^{s_{0}}\beta_{i}g(Z_{i})\big{]}\big{|}$ $=\big{|}\mathbf{E}\big{[}\Delta\widehat{W}\sum_{i=1}^{s_{0}}\beta_{i}(g(Z_{i})-g(y))\big{]}\big{|}$
$\leq K(1+|y|)\mathbf{E}(h|\Delta\widehat{W}|+|\Delta\widehat{W}|^{2})\leq K(1+|y|)h.$
In addition, by Lemma 3.1, $\big{|}\mathbf{E}\big{[}h\sum_{i=1}^{s_{0}}\alpha_{i}f(Z_{i})\big{]}\big{|}\leq Kh\sum_{i=1}^{s_{0}}(1+\mathbf{E}|Z_{i}|)\leq K(1+|y|)h$. Thus, we arrive at $|\mathbf{E}(Y^{SRK}_{t,y}(t+h)-y)|\leq K(1+|y|)h$. Finally, [16, Page 102, Lemma 2.2] and the Hölder inequality yield the desired result. ∎
Proof of Theorem 3.3
Proof.
In this proof, denote by $\iota$ the growing degree of a function of $\mathbf{F}$, which may vary for each appearance. It follows from the Taylor formula and $\mathcal{D}^{2}f,\,\mathcal{D}^{3}g\in\mathbf{F}$ that $\mathcal{D}f,\,\mathcal{D}^{i}g\in\mathbf{F}$ for $i=1,2$. The Taylor expansion gives
$g(Z_{j})=g(y)+\int_{0}^{1}\nabla g(y+\lambda(Z_{j}-y))(Z_{j}-y)\mathrm{d}\lambda=:g(y)+R^{1}_{g,j}.$ (A.1)
Since $g\in\mathbf{C}^{1}(\mathbb{R}^{d})$ and $g$ is Lipschitz continuous, $\nabla g$ is bounded. Accordingly, by Lemma 3.1, $\|R^{1}_{g,j}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|)h^{1/2}$ for $p\geq 1$. Similarly, one has
$f(Z_{j})=f(y)+R^{1}_{f,j}$ (A.2)
with $\|R^{1}_{f,j}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|)h^{1/2}$. Note that (3.11) is equivalent to
$Z_{i}=y+h\sum_{j=1}^{s_{0}}a_{ij}f(Z_{j})+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(Z_{j}),~i=1,\ldots,s_{0}.$ (A.3)
Plugging (A.1) into (A.3) yields
$Z_{i}=y+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(y)+R^{1}_{Z_{i}},$ (A.4)
where $R^{1}_{Z_{i}}=h\sum_{j=1}^{s_{0}}a_{ij}f(Z_{j})+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}R^{1}_{g,j}$. It follows from (A.2) and the Hölder inequality that $\|R^{1}_{Z_{i}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|)h$.
It follows from (A.4) and the Taylor formula that
$g(Z_{i})$ $=g(y)+\nabla g(y)(Z_{i}-y)+\tilde{R}^{2}_{g,i}=g(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla g(y)g(y)+R^{2}_{g,i},$ (A.5)
where $\tilde{R}^{2}_{g,i}:=\int_{0}^{1}(1-\lambda)\mathcal{D}^{2}g(y+\lambda(Z_{i}-y))(Z_{i}-y,Z_{i}-y)\mathrm{d}\lambda$ and $R^{2}_{g,i}=\nabla g(y)R^{1}_{Z_{i}}+\tilde{R}^{2}_{g,i}$. By Lemma 3.1, the boundedness of $h+|\Delta\widehat{W}|$ on $(0,h_{1}]$ and $\mathcal{D}^{2}g\in\mathbf{F}$,
$\|\mathcal{D}^{2}g(y+\lambda(Z_{i}-y))\|_{\otimes}\leq K(1+|y|^{\iota}),$ (A.6)
which combined with Lemma 3.1 leads to $\|\tilde{R}^{2}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})(\mathbf{E}|Z_{i}-y|^{2p})^{1/p}\leq K(1+|y|^{\iota})h$. Hence, we arrive at $\|R^{2}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h$. Similarly, one can prove
$f(Z_{i})=f(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla f(y)g(y)+R^{2}_{f,i}$ (A.7)
with $\|R^{2}_{f,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h$. Plugging (A.2) and (A.5) into (A.3) gives
$Z_{i}=y+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(y)+h\sum_{j=1}^{s_{0}}a_{ij}f(y)+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla g(y)g(y)+R^{2}_{Z_{i}},$ (A.8)
where $R^{2}_{Z_{i}}=h\sum_{j=1}^{s_{0}}a_{ij}R^{1}_{f,j}+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}R^{2}_{g,j}$. The previous estimates for $R^{1}_{f,j}$ and $R^{2}_{g,j}$ and the Hölder inequality yield $\|R^{2}_{Z_{i}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$.
The Taylor formula gives
$g(Z_{i})=g(y)+\nabla g(y)(Z_{i}-y)+\frac{1}{2}\mathcal{D}^{2}g(y)(Z_{i}-y,Z_{i}-y)+\tilde{R}^{3}_{g,i},$
$\tilde{R}^{3}_{g,i}:=\frac{1}{2}\int_{0}^{1}(1-\lambda)^{2}\mathcal{D}^{3}g(y+\lambda(Z_{i}-y))(Z_{i}-y,Z_{i}-y,Z_{i}-y)\mathrm{d}\lambda$. Similar to (A.6), it holds that $\|\mathcal{D}^{3}g(y+\lambda(Z_{i}-y))\|_{\otimes}\leq K(1+|y|^{\iota})$. Then the assumptions on $g$ and Lemma 3.1 produce $\|\tilde{R}^{3}_{g,j}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$. Plugging (A.8) into $\nabla g(y)(Z_{i}-y)$ and (A.4) into $\frac{1}{2}\mathcal{D}^{2}g(y)(Z_{i}-y,Z_{i}-y)$, we obtain
$g(Z_{i})=g(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla g(y)g(y)+R^{3}_{g,i},$ (A.9)
where
$R^{3}_{g,i}:=$ $\;h\sum_{j=1}^{s_{0}}a_{ij}\nabla g(y)f(y)+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}(\nabla g(y))^{2}g(y)+\nabla g(y)R^{2}_{Z_{i}}$
$\;+\frac{1}{2}\Delta\widehat{W}^{2}(\sum_{j=1}^{s_{0}}b_{ij})^{2}\mathcal{D}^{2}g(y)(g(y),g(y))+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\mathcal{D}^{2}g(y)(g(y),R^{1}_{Z_{i}})$
$\;+\frac{1}{2}\mathcal{D}^{2}g(y)(R^{1}_{Z_{i}},R^{1}_{Z_{i}})+\tilde{R}^{3}_{g,i}.$ (A.10)
Combining the previous estimates for $R^{1}_{Z_{i}}$, $R^{2}_{Z_{i}}$ and $\tilde{R}^{3}_{g,i}$, one has
$\|{R}^{3}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h,~p\geq 1\quad\text{and}\quad|\mathbf{E}(\Delta\widehat{W}{R}^{3}_{g,i})|\leq K(1+|y|^{\iota})h^{2}.$
Plugging (A.7) and (A.9) into (3.10) and using $\alpha^{\top}e=\beta^{\top}e=1$, $\beta^{\top}(Be)=\frac{1}{2}$, we have
$Y^{SRK}_{t,y}(t+h)=y+\Delta\widehat{W}g(y)+\frac{1}{2}\Delta\widehat{W}^{2}\nabla g(y)g(y)+hf(y)+R^{1}_{Y^{SRK}},$ (A.11)
where $R^{1}_{Y^{SRK}}:=\Delta\widehat{W}\sum_{i=1}^{s_{0}}\beta_{i}R^{3}_{g,i}+h\Delta\widehat{W}\sum_{i=1}^{s_{0}}\alpha_{i}\sum_{j=1}^{s_{0}}b_{ij}\nabla f(y)g(y)+h\sum_{i=1}^{s_{0}}\alpha_{i}R^{2}_{f,i}$. In addition, it holds that
$|\mathbf{E}R^{1}_{Y^{SRK}}|\leq K(1+|y|^{\iota})h^{2},\quad\|R^{1}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}.$ (A.12)
Define the operator $L:\mathbf{C}^{2}(\mathbb{R}^{d})\to\mathbf{C}(\mathbb{R}^{d})$ by $(Lu)(x)=\mathcal{D}u(x)(f(x)+\frac{1}{2}\nabla g(y)g(y))+\frac{1}{2}\text{trace}(\mathcal{D}^{2}u(x)g(x)g(x)^{\top})$, which is the generator of (3.2). It follows from the Itô formula that
$\;Y_{t,y}(t+h)$
$=$ $\;y+\int_{t}^{t+h}f(Y_{t,y}(s))\mathrm{d}s+\frac{1}{2}\int_{t}^{t+h}\nabla g(Y_{t,y}(s))g(Y_{t,y}(s))\mathrm{d}s+\int_{t}^{t+h}g(Y_{t,y}(s))\mathrm{d}W(s)$
$=$ $\;y+f(y)h+\int_{t}^{t+h}\int_{t}^{s}Lf(Y_{t,y}(r))\mathrm{d}r\mathrm{d}s+\int_{t}^{t+h}\int_{t}^{s}\nabla f(Y_{t,y}(r))g(Y_{t,y}(r))\mathrm{d}W(r)\mathrm{d}s$
$\;+\frac{1}{2}\nabla g(y)g(y)h+\frac{1}{2}\int_{t}^{t+h}\int_{t}^{s}L(\nabla gg)(Y_{t,y}(r))\mathrm{d}r\mathrm{d}s+\frac{1}{2}\int_{t}^{t+h}\int_{t}^{s}\nabla(\nabla gg)(Y_{t,y}(r))g(Y_{t,y}(r))$
$\;\mathrm{d}W(r)\mathrm{d}s+g(y)\Delta W+\int_{t}^{t+h}\int_{t}^{s}Lg(Y_{t,y}(r))\mathrm{d}r\mathrm{d}W(s)+\int_{t}^{t+h}\int_{t}^{s}\nabla g(Y_{t,y}(r))g(Y_{t,y}(r))\mathrm{d}W(r)\mathrm{d}W(s)$
==:= : $\;y+\Delta Wg(y)+\frac{1}{2}\nabla g(y)g(y)h+hf(y)+\int_{t}^{t+h}\int_{t}^{s}\nabla g(Y_{t,y}(r))g(Y_{t,y}(r))\mathrm{d}W(r)\mathrm{d}W(s)+\tilde{R}^{1,1}_{Y}.$
For any $\varphi\in\mathbf{F}$ and $p\geq 1$, it holds that $\|\varphi(Y_{t,y}(t+h))\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})$ owing to a similar estimate as (3.3). Further, the BDG inequality and the Hölder inequality yield $|\mathbf{E}\tilde{R}^{1,1}_{Y}|\leq K(1+|y|^{\iota})h^{2}$ and $\|\tilde{R}^{1,1}_{Y}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$. Again by the Itô formula and $\int_{t}^{t+h}\int_{t}^{s}\mathrm{d}W(r)\mathrm{d}W(s)=\frac{1}{2}(\Delta W^{2}-h)$, we have that
$\int_{t}^{t+h}\int_{t}^{s}\nabla g(Y_{t,y}(r))g(Y_{t,y}(r))\mathrm{d}W(r)\mathrm{d}W(s)=\frac{1}{2}(\Delta W^{2}-h)\nabla g(y)g(y)+\tilde{R}^{1,2}_{Y}$
with $\mathbf{E}\tilde{R}^{1,2}_{Y}=0$ and $\|\tilde{R}^{1,2}_{Y}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$. Letting $R^{1}_{Y}=\tilde{R}^{1,1}_{Y}+\tilde{R}^{1,2}_{Y}$, we get
$Y_{t,y}(t+h)=y+\Delta Wg(y)+\frac{1}{2}\Delta W^{2}\nabla g(y)g(y)+hf(y)+R^{1}_{Y}$ (A.13)
with $|\mathbf{E}{R}^{1}_{Y}|\leq K(1+|y|^{\iota})h^{2}$ and $\|{R}^{1}_{Y}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$.
By (A.11) and (A.13),
$Y_{t,y}(t+h)-Y^{SRK}_{t,y}(t+h)=(\Delta W-\Delta\widehat{W})g(y)+\frac{1}{2}\nabla g(y)g(y)(\Delta W^{2}-\Delta\widehat{W}^{2})+R^{1}_{Y}-R^{1}_{Y^{SRK}}.$
By (3.8), we have that for any $\varepsilon>0$, $|\mathbf{E}(\Delta W^{2}-\Delta\widehat{W}^{2}|=h|\mathbf{E}(\xi^{2}-\zeta_{h}^{2})|\leq K(\varepsilon)h^{\kappa+1-\varepsilon}$. From (3.6) we derive that $(\mathbf{E}|\Delta W-\Delta\widehat{W}|^{2p})^{1/(2p)}\leq K(\varepsilon)h^{\frac{1}{2}+\frac{\kappa}{2p}}$ for $p\geq 1$. Using (3.7) gives that $(\mathbf{E}|\Delta W^{2}-\Delta\widehat{W}^{2}|^{2p})^{1/(2p)}\leq K(\varepsilon)h^{1+\frac{\kappa}{2p}-\varepsilon}$. Therefore, it follows from the previous estimates for $R^{1}_{Y^{SRK}}$ and $R^{1}_{Y}$ and $\kappa\geq 2$ that for any $p\in[1,\frac{\kappa}{2}]$,
$|\mathbf{E}(Y_{t,y}(t+h)-Y^{SRK}_{t,y}(t+h))|$ $\leq K(1+|y|^{\iota})h^{2},$
$\|Y_{t,y}(t+h)-Y^{SRK}_{t,y}(t+h)\|_{\mathbf{L}^{2p}(\Omega)}$ $\leq K(1+|y|^{\iota})h^{3/2}.$
This verifies the condition (A1) of Theorem 2.1. Note that the condition (A3) of Theorem 2.1 holds for $Y_{t,y}(t+h)$ (see [21, Lemma 2.2]). In addition, condition (A2) of Theorem 2.1 is fulfilled owing to Lemma 3.2 and (3.3). Thus, the proof is completed as a result of Theorem 2.1. ∎
Proof of Lemma 3.4
Proof.
In this proof, denote by $\iota$ the growing degree of a function of $\mathbf{F}$, which may vary for each appearance. Note that $\mathcal{D}^{i}f\in\mathbf{F}$ for $i=1,2$ and $\mathcal{D}^{i}g\in\mathbf{F}$ for $i=1,2,3$ due to $\mathcal{D}^{3}f\in\mathbf{F}$ and $\mathcal{D}^{4}g\in\mathbf{F}$.
By (A.9)-(A.10), we can write
$g(Z_{i})=$ $\;g(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla g(y)g(y)+h\sum_{j=1}^{s_{0}}a_{ij}\nabla g(y)f(y)+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}(\nabla g(y))^{2}g(y)$
$\;+\frac{1}{2}\Delta\widehat{W}^{2}(\sum_{j=1}^{s_{0}}b_{ij})^{2}\mathcal{D}^{2}g(y)(g(y),g(y))+\bar{R}^{3}_{g,i},,$ (A.14)
with $\|\bar{R}^{3}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{3/2}$. Plugging (A.7) and (A.14) into (A.3) produces
$Z_{i}=$ $\;y+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(y)+h\sum_{j=1}^{s_{0}}a_{ij}f(y)+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla g(y)g(y)$
$\;+\Delta\widehat{W}h\sum_{j=1}^{s_{0}}a_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla f(y)g(y)+\Delta\widehat{W}h\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}a_{jk}\nabla g(y)f(y)$
$\;+\Delta\widehat{W}^{3}\sum_{j,k,l=1}^{s_{0}}b_{ij}b_{jk}b_{kl}(\nabla g(y))^{2}g(y)+\frac{1}{2}\Delta\widehat{W}^{3}\sum_{j=1}^{s_{0}}b_{ij}\big{(}\sum_{k=1}^{s_{0}}b_{jk}\big{)}^{2}\mathcal{D}^{2}g(y)(g(y),g(y))+R^{3}_{Z_{i}},$ (A.15)
where $R^{3}_{Z_{i}}:=h\sum_{j=1}^{s_{0}}a_{ij}R^{2}_{f,j}+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\bar{R}^{3}_{g,j}$. Since $\|{R}^{2}_{f,j}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h$ (see the proof of Theorem 3.3), we have $\|{R}^{3}_{Z_{i}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{2}$.
The Taylor formula leads to
$f(Z_{i})=f(y)+\nabla f(y)(Z_{i}-y)+\frac{1}{2}\mathcal{D}^{2}f(y)(Z_{i}-y,Z_{i}-y)+\tilde{R}^{3}_{f,i}$
with $\tilde{R}^{3}_{f,i}:=\frac{1}{2}\int_{0}^{1}(1-\lambda)^{2}\mathcal{D}^{3}f(y+\lambda(Z_{i}-y))(Z_{i}-y,Z_{i}-y,Z_{i}-y)\mathrm{d}\lambda$. By Lemma 3.1 and $\mathcal{D}^{3}f\in\mathbf{F}$,
$\|\tilde{R}^{3}_{f,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{3/2}.$ (A.16)
Plugging (A.8) into $\nabla f(y)(Z_{i}-y)$ and (A.4) into $\frac{1}{2}\mathcal{D}^{2}f(y)(Z_{i}-y,Z_{i}-y)$ yields
$\;f(Z_{i})=f(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla f(y)g(y)+h\sum_{j=1}^{s_{0}}a_{ij}\nabla f(y)f(y)$
$\;+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla f(y)\nabla g(y)g(y)+\frac{1}{2}\Delta\widehat{W}^{2}(\sum_{j=1}^{s_{0}}b_{ij})^{2}\mathcal{D}^{2}f(y)(g(y),g(y))+R^{3}_{f,i},$ (A.17)
where $R^{3}_{f,i}:=\nabla f(y)R^{2}_{Z_{i}}+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\mathcal{D}^{2}f(y)(g(y),R^{1}_{Z_{i}})+\frac{1}{2}\mathcal{D}^{2}f(y)(R^{1}_{Z_{i}},R^{1}_{Z_{i}})+\tilde{R}^{3}_{f,i}$. Combining the estimates for $R^{1}_{Z_{i}}$ and $R^{2}_{Z_{i}}$ from the proof of Theorem 3.3 and using (A.16), one has $\|R^{3}_{f,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{3/2}.$
Applying the Taylor formula to $g(Z_{i})$, one has
$g(Z_{i})=g(y)+\nabla g(y)(Z_{i}-y)+\frac{1}{2}\mathcal{D}^{2}g(y)(Z_{i}-y,Z_{i}-y)+\frac{1}{6}\mathcal{D}^{3}g(y)(Z_{i}-y,Z_{i}-y,Z_{i}-y)+\tilde{R}^{4}_{g,i},$
where $\tilde{R}^{4}_{g,i}:=\frac{1}{6}\int_{0}^{1}(1-\lambda)^{3}\mathcal{D}^{4}g(y+\lambda(Z_{i}-y))(Z_{i}-y,Z_{i}-y,Z_{i}-y,Z_{i}-y)\mathrm{d}\lambda$. By $\mathcal{D}^{4}g\in\mathbf{F}$ and Lemma 3.1, $\|\tilde{R}^{4}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{2}$. Plugging (A.15), (A.8) and (A.4) into $\nabla g(y)(Z_{i}-y)$, $\mathcal{D}^{2}g(y)(Z_{i}-y,Z_{i}-y)$ and $\mathcal{D}^{3}g(y)(Z_{i}-y,Z_{i}-y,Z_{i}-y)$, respectively, and keeping those terms with strong order no larger than 1.51.51.5, we obtain
$\;g(Z_{i})=g(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla g(y)g(y)+h\sum_{j=1}^{s_{0}}a_{ij}\nabla g(y)f(y)+\Delta\widehat{W}^{2}\Big{[}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}(\nabla g(y))^{2}g(y)$
$\;+\frac{1}{2}(\sum_{j=1}^{s_{0}}b_{ij})^{2}\mathcal{D}^{2}g(y)(g(y),g(y))\Big{]}+\Delta\widehat{W}h\Big{[}\sum_{j=1}^{s_{0}}a_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla g(y)\nabla f(y)g(y)$
$\;+\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}a_{jk}(\nabla g(y))^{2}f(y)+(\sum_{j=1}^{s_{0}}a_{ij})(\sum_{j=1}^{s_{0}}b_{ij})\mathcal{D}^{2}g(y)(f(y),g(y))\Big{]}$
$\;+\Delta\widehat{W}^{3}\Big{[}\sum_{j,k,l=1}^{s_{0}}b_{ij}b_{jk}b_{kl}(\nabla g(y))^{3}g(y)+\frac{1}{2}\sum_{j=1}^{s_{0}}b_{ij}(\sum_{k=1}^{s_{0}}b_{jk})^{2}\nabla g(y)\mathcal{D}^{2}g(y)(g(y),g(y))$
$\;+\sum_{j=1}^{s_{0}}b_{ij}\sum_{k,l=1}^{s_{0}}b_{ik}b_{kl}\mathcal{D}^{2}g(y)(g(y),\nabla g(y)g(y))+\frac{1}{6}(\sum_{j=1}^{s_{0}}b_{ij})^{3}\mathcal{D}^{3}g(y)(g(y),g(y),g(y))\Big{]}+R^{4}_{g,i}$ (A.18)
with $\|R^{4}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{2}$.
Plugging (Proof of Lemma 3.4)-(A.18) into (3.10) and using $\alpha^{\top}e=\beta^{\top}e=1$ and $\beta^{\top}(Be)=\frac{1}{2}$, we obtain (3.12) with $R^{2}_{Y^{SRK}}=h\sum_{i=1}^{s_{0}}\alpha_{i}R^{3}_{f,i}+\Delta\widehat{W}\sum_{i=1}^{s_{0}}\beta_{i}R^{4}_{g,i}$. Finally, $\|R^{2}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\alpha^{\prime}})h^{5/2}$ is obtained due to the previous estimates for $R^{3}_{f,i}$ and $R^{4}_{g,i}$. ∎